### v1

# Hubbard and Heisenberg models on hyperbolic lattices - Metal-insulator transitions, global antiferromagnetism and enhanced boundary fluctuations

Abstract
We study the Hubbard and Heisenberg models on hyperbolic lattices with open boundary conditions by means of mean-field approximations, spin-wave theory and quantum Monte Carlo (QMC) simulations. For the Hubbard model we use the auxiliary-field approach and for Heisenberg systems the stochastic series expansion algorithm. We concentrate on bipartite lattices where the QMC simulations are free of the negative sign problem. The considered lattices are characterized by a Dirac like density of states, Schläfli indices $\{p,q\}=\{10,3\}$ and $\{8,3\}$, as well as by flat bands, $\{8,8\}$. The Dirac density of states cuts off the logarithmic divergence of the staggered spin susceptibility and allows for a finite $U$ semi-metal to insulator transition. This transition has the same mean-field exponents as for the Euclidean counterpart. In the presence of flat bands we observe the onset of magnetic ordering at any finite $U$. The magnetic state at intermediate coupling can be described as a global antiferromagnet. It breaks the $C_{p}$ rotational and time reversal symmetries but remains invariant under combined $C_{p}\mathcal{T}$ transformations. The state is characterized by macroscopic ferromagnetic moments, that globally cancel. We observe that fluctuations on the boundary of the system are greatly enhanced: while spin wave calculations predict the breakdown of antiferromagnetism on the boundary but not in the bulk, QMC simulations show a marked reduction of the staggered moment on the edge of the system.
I Introduction
The effect of dimensionality and point group symmetries on correlation induced phenomena is an established research domain in flat Euclidean space. The aim of this paper is to investigate the effect curvature has on the physics of standard models of correlated electrons and spins. The role of the curvature is widely studied in models of statistical mechanics, like the Ising model [1, 2, 3, 4, 5, 6, 7, 8]. A constant negative curvature can for example lead to the presence of symmetry-broken phases and phase transitions, not present in the flat Euclidean plane [1, 9, 2, 10, 11] or change the critical properties of phase transitions [12, 3, 5]. Also in condensed matter physics the popularity of hyperbolic lattices is increasing. The topics of interest are both single-particle physics, like the electronic band structure [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] – also under the influence of a magnetic field and topological phenomena – and correlation effects [28, 29, 26]. The motivation can partially be attributed to recent experimental realizations of hyperbolic lattices in circuit quantum electrodynamics and electric circuits [30, 31].
In this paper we study the Hubbard and Heisenberg models on three exemplary regular tessellations of the hyperbolic plane, that exhibit different electronic properties in the non-interacting limit. We use a Hartree-Fock approximation, a spin wave analysis, auxiliary-field and stochastic series expansion quantum Monte Carlo (QMC) methods to investigate the models as a function of the interaction strength and system size. Since all considered lattices are bipartite, we can perform QMC simulations without a negative sign problem for both the Hubbard and Heisenberg models.
The electronic properties of the three lattices resemble a Dirac system with a linear density of states (DOS) around the Fermi level ($\{10,3\}$ and $\{8,3\}$ lattice), and a flat-band system with a high DOS at the Fermi level ($\{8,8\}$ lattice). The mean-field approximation of the Hubbard model on the $\{10,3\}$ lattice is already studied in Ref. [26], and we complement these results with both spin wave calculations and first-time fermionic QMC studies on the hyperbolic lattice. We will compare the spatial distribution of the magnetic ordering in the weak-coupling and strong-coupling limits, the transition from a semi-metal to insulator with antiferromagnetic (AFM) ordering as a function of the interaction strength and thermodynamic properties.
The remainder of the paper is organized as follows. In Sec. II we define the Hubbard and Heisenberg models and the hyperbolic lattices on which we study them. In Sec. III we introduce the mean-field approximation, spin-wave analysis and the QMC algorithms that will be used in this work. In Sec. IV we present our numerical results before we conclude in Sec. V. In the Appendix additional numerical data can be found.
II Hyperbolic Lattices, Hubbard and Heisenberg Models
II.1 Hyperbolic lattices
Figure 1: $\{8,3\}$ lattice in the Poincaré disk model with increasing number of layers $n_{l}$. Reference angle marked for latter use.
In this work, we consider two-dimensional graphs that can be categorized by their Schläfli index $\{p,q\}$. The lattice sites are located at the corners of regular $p$-gons and every lattice site has coordination number $q$ [32]. Hyperbolic lattices with a constant negative curvature fulfill the condition $(p-2)(q-2)>4$. In contrast, lattices with $(p-2)(q-2)=4$ lie within the flat Euclidean plane and lattices with $(p-2)(q-2)<4$ in a positively curved plane. For the Euclidean plane exactly three solutions exist – the square lattice $\{4,4\}$, the honeycomb lattice $\{6,3\}$ and the triangular lattice $\{3,6\}$, whereas for the hyperbolic case infinitely many pairs $\{p,q\}$ exist that fulfill the condition [33].
We concentrate on finite size hyperbolic lattices with open boundary conditions. The lattice size can be increased by starting with a single $p$-gon and successively adding shells or layers of $p$-gons. The number of lattice sites grows exponentially with the number of layers, strongly limiting the accessible lattice sizes for our calculations. In Fig. 1 the $\{8,3\}$ lattice is shown with up to $n_{l}=3$ layers. In order to properly depict the lattices, the Poincaré disk model is used. All lattice sites are projected into the open unit disk and all edges, connecting two nearest-neighboring sites, are geodesic lines. Those geodesics are circular arcs in the Poincaré disk model, that would hit the unit circle perpendicular and they all have the same length within the hyperbolic metric  [^1] The metric in the hyperbolic plane with a constant negative curvature of -1 is given by $ds^{2}=4\frac{dx^{2}+dy^{2}}{(1-x^{2}-y^{2})^{2}}$. Here, $x$ and $y$ are the coordinates within the open unit circle in the Poincaré disk model. [35, 36].
Drawing conclusions from the results on finite lattices about the properties in the thermodynamic limit is somewhat more complicated than in the Euclidean plane. For open boundary conditions the ratio of lattice sites at the edge to the total number of lattice sites converges to a finite value, so the behavior of the edge sites is not negligible even for large system sizes [3, 6, 24, 27]. If the bulk properties are of interest, the lattice can be successively increased and the calculation can only be done for the inner bulk part, neglecting the outer shells [3, 4]. Another possibility is to use periodic boundary conditions. Some classes of hyperbolic tilings can be placed on compact surfaces with higher genus $g\geq 2$ with periodic boundary conditions [37, 8, 28, 18, 19]. But in this work we use open boundary conditions and also take into account the boundary contributions.
Fig. 1 also shows, that the $\{8,3\}$ lattice is bipartite with sublattice $A$ and $B$, like all lattices in this work. In general for even $p$ the resulting lattices are bipartite [27]. In Table 1 of the Appendix we provide a list of the considered lattice sizes and their respective number of lattice sites $N_{s}$.
II.2 Hubbard and Heisenberg models
The Hamiltonian of the Hubbard model is given by [38, 39, 40]
$\hat{H}$ $=$ $\hat{H}_{t}+\hat{H}_{U}\,,$
$\hat{H}_{t}$ $=$ $-t\sum_{\langle\bm{i},\bm{j}\rangle}\sum_{\sigma=1}^{N_{\sigma}}\left(\hat{c}_{\bm{i},\sigma}^{\dagger}\hat{c}_{\bm{j},\sigma}^{\phantom{\dagger}}+\mathrm{h.c.}\right)\,,$
$\hat{H}_{U}$ $=$ $\frac{U}{N_{\sigma}}\sum_{\bm{i}}\left(\sum_{\sigma=1}^{N_{\sigma}}\hat{n}_{\bm{i},\sigma}-\frac{N_{\sigma}}{2}\right)^{2}\,.$ (1)
The first term describes the hopping of fermions on bonds $b=\langle\bm{i},\bm{j}\rangle$ connecting nearest-neighboring sites $\bm{i}$ and $\bm{j}$ with hopping strength $t$. The operator $\hat{c}_{\bm{i},\sigma}^{\dagger}$ creates a fermion in a Wannier state centered at site $\bm{i}$ and with $z$ component of spin equal to $\sigma$. We will consider $N_{\sigma}=2$ (spin $S=\frac{1}{2}$) flavored fermions. The fermions interact via a repulsive on-site Hubbard potential with strength $U$, as described in the second term.
The model is symmetric under global SU(2) rotations with the spin operators as the generators of the group
$\hat{S}^{\alpha}=\frac{1}{2}\sum_{\bm{i},\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{\bm{i},\sigma}\sigma^{\alpha}_{\sigma,\sigma^{\prime}}\hat{c}^{\phantom{\dagger}}_{\bm{i},\sigma^{\prime}}\,.$ (2)
Here $\sigma^{\alpha=x,y,z}$ are the three Pauli matrices. Furthermore, the model is invariant under particle-hole transformations $\hat{P}^{-1}\hat{c}^{\dagger}_{\bm{i},\sigma}\hat{P}=(-1)^{\bm{i}}\hat{c}_{\bm{i},\sigma}^{\phantom{\dagger}}$ on a bipartite lattice at half-filling. The prefactor $(-1)^{\bm{i}}$ is a multiplicative phase factor, that takes the value $+1$ on sublattice $A$ and $-1$ on sublattice $B$.
The lattice geometry can significantly influence the electronic band structure and therefore the behavior of the electrons, when exposed to interactions. This effect can be observed for the available regular bipartite graphs in the Euclidean plane, the square lattice $\{4,4\}$ and the honeycomb lattice $\{6,3\}$. On both lattices the repulsive Hubbard model at half-filling undergoes a transition from a (semi-)metal to an AFM insulator in the limit of strong on-site interactions. The Fermi surface of the half-filled square lattice exhibits van Hove singularities leading to a logarithmically divergent DOS at the Fermi surface. This results in an essential singularity of the staggered magnetization in the Hartree-Fock approximation, meaning that for any finite interaction strength $U$ the AFM state is expected to have a lower energy than the paramagnetic solution [41, 42]. Numerical studies support the presence of long-range AFM order for any finite interaction $U$ [41, 43, 44, 45].
On the honeycomb lattice the band structure of the non-interacting case exhibits two Dirac cones and the Fermi surface is concentrated on two points with a vanishing DOS. A finite critical on-site repulsion $U_{c}$ is needed for the formation of long-range AFM ordering [46]. Numerical studies locate a continuous and direct semi-metal to insulator transition at $U_{c}/t\approx 3.8$ [47, 48, 49]. This transition belongs to the Gross-Neveu universality class [50, 48, 49].
For the hyperbolic case we know, that the unique ground state of the half-filled repulsive Hubbard model is a $S_{\mathrm{tot}}=0$ state for any bipartite lattice by a theorem of Lieb [51]. $S_{\mathrm{tot}}$ is the total spin and the theorem was proven for any bipartite collection of sites, meaning no explicit periodicity or dimensionality is assumed in the proof of the theorem. Furthermore in the strong-coupling limit the model maps onto the Heisenberg model with an AFM coupling [52], since the mapping is independent of the lattice. The specific form of the model reads:
$\hat{H}=J\sum_{\langle\bm{i},\bm{j}\rangle}\hat{\bm{S}}_{\bm{i}}\cdot\hat{\bm{S}}_{\bm{j}}$ (3)
with $J=\frac{4t^{2}}{U}$ and $\hat{\bm{S}}_{\bm{i}}=\frac{1}{2}\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{\bm{i},\sigma}\bm{\sigma}_{\sigma,\sigma^{\prime}}\hat{c}^{\phantom{\dagger}}_{\bm{i},\sigma^{\prime}}$
III Methods
III.1 Mean-field approximation
As a first approach to study the Hubbard model on hyperbolic lattices we choose a mean-field approximation. We decouple the electron-electron interaction by coupling the $z$ component of the spin operator to the fields $\phi_{\bm{i}}$
$\hat{H}_{\mathrm{MF}}$ $=$ $\hat{H}_{t}-U\sum_{\bm{i}}\phi_{\bm{i}}\left(\hat{n}_{\bm{i},\uparrow}-\hat{n}_{\bm{i},\downarrow}\right)+\frac{U}{2}\sum_{\bm{i}}\phi_{\bm{i}}^{2}\,,$ (4)
$\phi_{\bm{i}}$ $=$ $\langle\hat{n}_{\bm{i},\uparrow}-\hat{n}_{\bm{i},\downarrow}\rangle_{\phi}=2\langle\hat{S}_{\bm{i}}^{z}\rangle_{\phi}=m^{z}_{\bm{i}}\,,$ (5)
where $\langle\dots\rangle_{\phi}=\frac{\mathop{\mathrm{Tr}}[\dots e^{-\beta\hat{H}_{\mathrm{MF}}}]}{\mathop{\mathrm{Tr}}[e^{-\beta\hat{H}_{\mathrm{MF}}}]}$ is the expectation value with respect to $\hat{H}_{\mathrm{MF}}$ and $\beta=T^{-1}$ the inverse temperature. The fields $\phi_{\bm{i}}$ are equivalent to the local magnetic order parameter and due to the absence of translation symmetry in the considered lattices, we allow the fields to vary from site to site. In the following, Eq. (5) will be solved self-consistently. In Appendix C we show that, due to particle-hole symmetry, the staggered spin susceptibility diverges logarithmically in the low-temperature limit, the prefactor being the DOS at the Fermi energy. Hence, we foresee AFM ordering and define the staggered order parameter
$m^{z}=\frac{1}{N_{s}}\sum_{\bm{i}}(-1)^{\bm{i}}m^{z}_{\bm{i}}\,.$ (6)
III.2 Spin wave approximation
We continue with a spin wave approximation. In the strong-coupling limit the Hubbard model maps onto the AFM Heisenberg model independently of the lattice geometry [52]
$\hat{H}=\sum_{\langle\bm{i},\bm{j}\rangle}\hat{\bm{S}}_{\bm{i}}\cdot\hat{\bm{S}}_{\bm{j}}=\frac{1}{2}\sum_{\bm{i},\bm{j}}T_{\bm{i},\bm{j}}\hat{\bm{S}}_{\bm{i}}\cdot\hat{\bm{S}}_{\bm{j}}\,.$ (7)
The matrix $T$ is the adjacency matrix of the lattice with $T_{\bm{i},\bm{j}}=1$, if sites $\bm{i}$ and $\bm{j}$ are nearest neighbors and $T_{\bm{i},\bm{j}}=0$ otherwise. Following Ref. [53], we use a Holstein-Primakoff transformation to map the spin operators $\hat{\bm{S}}_{\bm{i}}$ to boson operators $\hat{b}_{\bm{i}}^{\dagger}$ and expand in $\frac{1}{S}$. Since we start from an AFM, we choose on sublattice $A$ the representation
$\hat{S}_{\bm{i}}^{z}=S-\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\,,\quad\hat{S}_{\bm{i}}^{+}=\hat{S}_{\bm{i}}^{x}+i\hat{S}_{\bm{i}}^{y}=\sqrt{2S}\hat{b}_{\bm{i}}^{\phantom{\dagger}}$ (8)
and on sublattice $B$
$\hat{S}_{\bm{i}}^{z}=\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}-S\,,\quad\hat{S}_{\bm{i}}^{+}=\sqrt{2S}\hat{b}_{\bm{i}}^{\dagger}\,.$ (9)
The correction to the classically expected Néel state is given by the expectation value of the bosonic occupation number $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$. In Appendix B we sketch the calculation of the site-dependent correction.
III.3 Quantum Monte Carlo
In order to acquire deeper information about the electronic correlations on the various lattice geometries and support the mean-field and spin wave data, we employ QMC studies as our secondary approach. In the weak-coupling limit, we use a finite-temperature version of the auxiliary-field QMC (AFQMC) method to simulate the Hubbard model [54, 44, 55]. For the strong-coupling regime, the Hubbard model is mapped to an AFM Heisenberg model, and in this regime we use the Stochastic Series Expansion (SSE) method with directed loop updates [56, 57].
III.3.1 Auxiliary-field QMC
We use the Algorithms for Lattice Fermions (ALF) [58] implementation of the finite-temperature AFQMC algorithm [54, 44, 55]. It allows us to accurately simulate the model in various parameter regimes. As for the Hubbard model at half-filling, our simulations do not suffer from the negative sign problem.
Although the construction of momentum space, similar to Euclidean lattices, is possible [16, 18, 19], it does not apply to our lattices with open boundary conditions due to the absence of translation invariance and we have no notion of momentum space. Yet, the QMC algorithm can be employed to calculate the uniform susceptibility
$\chi_{O}=\frac{\beta}{N_{s}}\sum_{\bm{i},\bm{j}}^{N_{s}}\left(\big{\langle}\hat{O}_{\bm{i}}\hat{O}_{\bm{j}}\big{\rangle}-\big{\langle}\hat{O}_{\bm{i}}\big{\rangle}\big{\langle}\hat{O}_{\bm{j}}\big{\rangle}\right)$ (10)
at nonzero Hubbard-$U$ with $\hat{O}_{\bm{i}}$ being either density $\hat{n}_{\bm{i}}=\sum_{\sigma}\hat{n}_{\bm{i},\sigma}$ or spin operators $\hat{S}_{\bm{i}}^{z}$. To directly compare the QMC results with mean-field, we make use of the bipartite nature of the chosen lattices by constructing an AFM order parameter as
$m_{\rm{AFM}}^{z}=\sqrt{\frac{1}{N_{s}^{2}}\sum_{\bm{i},\bm{j}}^{N_{s}}(-1)^{\bm{i}}(-1)^{\bm{j}}\big{\langle}\hat{S}_{\bm{i}}^{z}\hat{S}_{\bm{j}}^{z}\big{\rangle}}\,,$ (11)
which is directly related to the spin-spin-correlation function. For all of the presented data, we used a code for the Hubbard model with SU(2)-decoupling and an imaginary time discretization of $\Delta\tau t=0.1$ [58].
III.3.2 Stochastic Series Expansion
As a second QMC approach, we simulate the Heisenberg model on the hyperbolic lattices using the SSE method with directed loop updates [56, 57]. Since the considered lattices are of bipartite nature, the method can be formulated without a sign problem. This approach allows us to access larger lattice sizes compared to the AFQMC method, since SSE simulations scale as $\mathcal{O}(\beta N_{s})$ as opposed to $\mathcal{O}(\beta N_{s}^{3})$ [58].
The SSE method is based on the Taylor expansion of the partition function $Z$ in some complete basis $\{\left|\alpha\right>\}$ (here we choose the $S^{z}$-basis)
$Z=\sum_{\alpha}\sum_{n=0}^{\infty}\sum_{S_{n}}\frac{(-\beta)^{n}}{n!}\left<\alpha\right|\prod_{i=1}^{n}H_{a_{i},b_{i}}\left|\alpha\right>\,,$ (12)
$n$ is the current expansion order, $S_{n}$ specifies the operator string $[(a_{1},b_{1}),\ldots,(a_{n},b_{n})]$ and $H_{a_{i},b_{i}}$ is a bond term of the Hamiltonian which operates on bond $b_{i}$. This action can be diagonal or off-diagonal depending on $a_{i}$. The configuration space then consists of operator strings of varying size $n$. To sample this space, we use two main types of updates. Diagonal updates are a local type of update, in which we insert/remove a Hamiltonian operator in $S_{n}$. The directed loop update [57, 59] is a global type of update in which a loop is constructed and flipped in the vertex representation of $S_{n}$, allowing us to change a large part of the configuration in one step.
Both the uniform spin susceptibility [Eq. (10)] and the AFM order parameter [Eq. (11)] rely on the calculation of equal-time two-point spin correlations $\big{\langle}\hat{S}^{z}_{\bm{i}}\hat{S}^{z}_{\bm{j}}\big{\rangle}$. As these operators are diagonal in the $S^{z}$-basis, their evaluation within the SSE formalism is trivial.
Figure 2: (Ia)-(Ic) DOS $\rho(E)$ as a function of the energy $E$ at $\delta=0.01$ for the largest considered lattice sizes ($n_{l}=5$ for the $\{10,3\}$ lattice, $n_{l}=6$ for $\{8,3\}$ and $n_{l}=4$ for $\{8,8\}$, see also Table 1) and (IIa)-(IIc) uniform spin susceptibility $\chi_{S}$ as a function of the temperature $T$ and the number of layers $n_{l}$ in the non-interacting case $U=0$.
IV Results
In the following we study the SU(2) spin symmetric Hubbard and Heisenberg models on three different hyperbolic lattice geometries at the particle-hole symmetric point. As shown in Appendix C particle-hole symmetry leads to a logarithmic divergence of the staggered spin susceptibility, provided that the DOS is finite at the Fermi energy. Hence, the first question that we will address in Sec. IV.1 is the DOS in the non-interacting limit. In Sec. IV.2 we present and discuss our mean-field results. In Sec. IV.3, we then systematically take into account fluctuations with a spin wave approximation for the Heisenberg model, and numerically exact QMC simulations for both models.
IV.1 Non-interacting limit $U=0$
First we consider the DOS of the $\{10,3\}$ lattice,
$\rho(E)=-\frac{1}{\pi N_{s}}\sum_{n}\mathrm{Im}G^{R}(n,E)\,.$ (13)
The retarded Green function is given by $G^{R}(n,E)=\frac{1}{E-E_{n}+i\delta}$ with an infinitesimal $\delta$ and $E_{n}$ are the discrete energy eigenstates of the tight-binding Hamiltonian $\hat{H}_{t}$. Since our Hamiltonian enjoys particle-hole symmetry, the DOS is symmetric around $E=0$ and the chemical potential for half-band filling is pinned to zero for all temperatures. The uniform spin susceptibility $\chi_{S}$, defined in Eq. (10), is given by
[…]
for the free case. In the above, $f$ denotes the Fermi-function. Both the DOS and the uniform spin susceptibility are plotted in Fig. 2(Ia) and 2(IIa). The data, in the limit of large values of $n_{l}$, is consistent with a DOS $\rho(E)\propto|E|$ and a corresponding uniform spin susceptibility $\chi_{S}\propto T$. The DOS close to the Fermi level as well as the uniform spin susceptibility bear similarities with Dirac systems.
In Figs. 2(Ib) and 2(IIb) we consider the $\{8,3\}$ lattice. At even number of layers, there are no zero-energy eigenstates in the spectrum and $\chi_{S}$ decreases exponentially in the low-temperature limit. The activation gap can be attributed to the finite-size gap. On the other hand, for an odd number of layers exactly two eigenstates at zero energy are present, thereby resulting in a $T^{-1}$ law of the uniform spin susceptibility in the low-temperature limit. But as the number of layers is increased the ratio of zero-energy eigenstates to the total number of eigenstates quickly converges to zero and the magnitude of the $T^{-1}$ law vanishes in the large $n_{l}$ limit. We note that the slope of the DOS around $E=0$ and of the spin susceptibility is steeper for the $\{8,3\}$ lattice compared to the former lattice. In Euclidean space, the slope is given by the inverse of the Fermi velocity.
The $\{8,8\}$ lattice on the other hand shows flat band features on the considered finite lattices. A large number of energy eigenstates are located at zero energy [Fig. 2(Ic)] and the uniform spin susceptibility diverges in the zero temperature limit [Fig. 2(IIc)]. As mentioned previously, the boundary of hyperbolic lattices has an extensive number of sites such that the bulk, Ref. [27], and total DOS, Eq. (13) can differ. Comparing results, we see that the flat band observed in the $\{8,8\}$ lattice is boundary induced.
IV.2 Mean-field approximation
IV.2.1 Staggered magnetization and metal-insulator transition
Figure 3: (a)-(c) AFM order parameter $m^{z}$ as a function of the interaction strength $U$. The red line in (a) is a linear fit to the $n_{l}=4$ data of the $\{10,3\}$ lattice close to the critical point $U_{c}\approx 0.76t$ and in (b) to the $n_{l}=6$ data close to $U_{c}\approx 0.54t$. Insets: local order parameter $m_{\bm{i}}^{z}$ at $U=5.0t$ for (a) $n_{l}=4$, (b) $n_{l}=5$ and (c) $n_{l}=4$.
Figure 4: (a) AFM order parameter $m^{z}$ as a function of the temperature $T$ for different interaction strengths as obtained from the mean-field approximation. (b) Critical temperature $T_{c}$ as a function of the Hubbard interaction $U$. For the $\{10,3\}$ and $\{8,8\}$ lattices we chose $n_{l}=3$ layers and for the $\{8,3\}$ lattice $n_{l}=4$ layers.
In this section we study the three lattice systems with the mean-field approximation, introduced in Sec. III.1. Figure 3 shows the order parameter as a function of the interaction strength $U$. In the strong-coupling limit a clear AFM state develops for all lattices with a vanishing total spin and a strong AFM moment $m^{z}$. The insets of Fig. 3 show the local magnetic moment $m_{\bm{i}}^{z}$, with a nearly spatially uniform modulus $|m_{\bm{i}}^{z}|$, deep in the AFM phase at $U=5t$.
The behavior of the three lattices in the weak-coupling limit differ strongly as a consequence of the DOS in the non-interacting case. In fact in Appendix C we show, that for a finite DOS at the Fermi level and under the assumption of partial particle-hole symmetry an instability towards AFM can be expected even in the absence of translational symmetry.
For the Dirac-like $\{10,3\}$ lattice the order parameter $m^{z}$ picks up a nonzero value only above a critical interaction strength $U_{c}\approx 0.76t$. Below $U_{c}$ the paramagnetic state is energetically favorable. Hartree-Fock mean-field theory for the flat honeycomb lattice predicts a linear increase of the order parameter above the critical interaction $m^{z}\propto(U-U_{c})$, i.e. a critical exponent of $\beta=1$ [46, 60, 61]. The order parameter of the $n_{l}=4$ lattice is well described by a linear fit close to the critical point [Fig. 3(a)].
The order parameter on the $\{8,3\}$ exhibits clear odd-even effects in $n_{l}$. Let us consider the $n_{l}=1$ lattice, that corresponds to the $p$-site ring with $p=8$. Since the rotational symmetry of the ring is shared by the $n_{l}>1$ lattices, a symmetry analysis will prove to be very useful. The ground state of the $p=8$ site ring is four-fold degenerate:
|±$|\pm,\sigma\rangle=\hat{c}^{\dagger}_{k=\pm\frac{\pi}{2},\sigma}|0\rangle\,.$ (15)
In the above $|0\rangle$ is the non-degenerate ground state of the 8-site ring occupied with six electrons, and $\hat{c}^{\dagger}_{k,\sigma}=\frac{1}{\sqrt{p}}\sum_{j=1}^{p}e^{ikj}\hat{c}^{\dagger}_{j,\sigma}$. Under time-reversal symmetry, $\mathcal{T}$, and $C_{8}$ rotations the states transform as
$\mathcal{T}|\pm,\sigma\rangle=\left(i\sigma^{y}\right)_{\sigma,\sigma^{\prime}}|\mp,\sigma^{\prime}\rangle$ (16)
and
$C_{8}|\pm,\sigma\rangle=e^{\pm i\frac{\pi}{2}}|\pm,\sigma\rangle\,.$ (17)
In the presence of AFM ordering, time-reversal and $C_{8}$ symmetries are individually broken, but the combined $C_{8}\mathcal{T}$ symmetry is conserved. Terms such as
$\Delta\sum_{\sigma}\sigma\left[|+,\sigma\rangle\langle-,\sigma|+|-,\sigma\rangle\langle+,\sigma|\right]$ (18)
are allowed and will lift the ground state degeneracy, yielding
$|\Psi_{0}\rangle=\frac{1}{2}\left(|+,\uparrow\rangle+|-,\uparrow\rangle\right)\otimes\left(|+,\downarrow\rangle-|-,\downarrow\rangle\right)\,.$ (19)
This state has a finite staggered magnetization $m^{z}$ for any value of $\Delta$, and is symmetric under combined $C_{8}\mathcal{T}$ symmetries. The jump in the magnetization observed at $n_{l}=1$ is a consequence of the degeneracy of the non-interacting ground state and the concomitant divergence of the uniform spin susceptibility.
While the symmetry analysis will apply to arbitrary values of $n_{l}$, the degeneracy only occurs for odd values of $n_{l}$. Since the degeneracy is intensive, the jump in the staggered magnetization will be suppressed as a function of $n_{l}$. For even values of $n_{l}$ the finite size gap leads to a finite value of $U$ beyond which magnetic ordering will occur.
In the large $n_{l}$ limit we conjecture a Dirac-like system. The–in comparison to the $\{10,3\}$ lattice lower–critical interaction strength, that lies within the range of $U_{c}\simeq 0.5t-0.6t$, is consistent with the observation of Sec. IV.1, that the slope of the DOS around $E=0$ and of the uniform spin susceptibility of the $\{8,3\}$ lattice is steeper than the one of the $\{10,3\}$ lattice. Mean-field analysis for a Dirac system in flat space predicts a linear proportionality $U_{c}\propto v_{\mathrm{F}}$ [61].
In Fig. 4(a) we present the AFM order parameter as a function of temperature $T$, and Fig. 4(b) the critical temperature as a function of the interaction strength. For the Dirac-like systems our data is consistent with $m^{z}\propto(T-T_{c})^{1/2}$, in the vicinity of $T_{c}$ and $T_{c}\propto(U-U_{c})$, in the vicinity of $U_{c}$. These forms are precisely the same as obtained for Dirac systems on Euclidean lattices. Hence at the mean-field level the semi-metal to insulator transition in hyperbolic and flat spaces are identical.
In contrast the flat-band-like $\{8,8\}$ lattice immediately supports AFM ordering for any finite $U$ due to the large DOS at the Fermi level in the non-interacting limit. The singularity at $U=0$ for all considered values of $n_{l}$ can be traced back to the divergence of the uniform spin susceptibility and accompanying extensive ground state degeneracy. Although the $T=0$ magnetization is singular at $U=0$ the Néel transition temperature seems to follow a powerlaw [Fig. 4)] as for other flat band systems [62]. In particular $\lim_{T\rightarrow 0}\lim_{U\rightarrow 0}m^{z}=0$, but $\lim_{U\rightarrow 0}\lim_{T\rightarrow 0}m^{z}>0$.
Figure 5: (a) Energy eigenvalues $E_{n}$ as obtained from the self-consistent solution of the mean-field Hamiltonian (4), inset: close-up of the Fermi level at $E_{n}=0$, (b) absolute value of the wave function of zero mode $|\psi_{n=1444}(\bm{i})|^{2}$ at $U=0.001t$ and $\sigma=\uparrow$ and (c) magnetization $m^{z}_{\bm{i}}$ at $U=0.4t$ for the $\{8,3\}$ lattice with $n_{l}=5$ layers.
Figure 6: Local magnetization $m^{z}_{\bm{i}}$ as obtained from mean-field approximation as a function of the angle $\phi$ at two different interaction strengths (a) $U=0.4t$ and (b) $U=5t$ for the $\{8,3\}$ lattice with $n_{l}=3$. For clarity all sites of one layer are in a separate plot [(I) layer 1, (II) layer 2 and (III) layer 3].
IV.2.2 Global antiferromagnetism
Using the $\{8,3\}$ lattice as an example, we now concentrate on the weak-coupling limit, meaning that the interaction is (much) smaller than the electronic bandwidth in the non-interacting limit. Fig. 5(a) shows the energy eigenvalues of the self-consistent solution of the mean-field Hamiltonian (4) for the $\{8,3\}$ lattice with $n_{l}=5$ layers and different interaction strengths $U$. In the non-interacting limit the spectrum is gapless with exactly two zero-energy eigenstates. As for the $n_{l}=1$ case, the wave function of a single zero mode, depicted in Fig. 5(b), shows broken $C_{8}$ symmetry. Turning on a finite interaction opens a gap, [inset of Fig. 5(a)] and the AFM order parameter acquires a finite value.
As for the $n_{l}=1$ case, a small interaction strength of e.g. $U=0.4t$ affects mainly the two eigenstates at the Fermi level [inset of Fig. 5(a)]. Figure 5(b) shows the absolute value of the wave function $|\psi_{n}(\bm{i})|^{2}$ of one zero mode in real-space. In order to lift the degeneracy of the zero modes and have a uniquely defined pattern of the wave function, we choose a small interaction strength of $U=0.001t$ in Fig. 5(b). The wave function has support only on a small number of sites and it breaks the $C_{p=8}$ symmetry of the lattice down to $C_{p/2=4}$. The wave function of the second zero mode is similar, but rotated by $\pi/4$.
The spatial distribution of the order parameter $m^{z}_{\bm{i}}$ follows the pattern of these wave functions [Fig. 5(c)]. In particular, at finite $U$ the ground state is unique, breaks time-reversal and $C_{8}$ symmetries, but satisfies:
$C_{8}\mathcal{T}|\Psi_{0}\rangle=|\Psi_{0}\rangle\,.$ (20)
For the $n_{l}=1$ case, this just results in AFM, since the $C_{8}$ transformation amounts to a unit translation along the chain. For $n_{l}>1$, new magnetic structures emerge, that we will identify with global AFM in these hyperbolic lattices as seen e.g. in strained graphene [63].
First of all, the total magnetic moment vanishes. However, the vanishing of this moment involves cancellation of extensive ferromagnetic moments on each $p$-th sector of the lattice. This is depicted in Fig. 6, where the local magnetization $m_{\bm{i}}^{z}$ is plotted as a function of the angle $\phi$ in the different layers, both in the weak $U=0.4t$ and strong-coupling limit $U=5t$ for the $\{8,3\}$ lattice with $n_{l}=3$ layers. In the strong-coupling limit the magnetization is almost uniform on the whole lattice. At weak coupling the magnetization varies strongly with the angle. Summing over all sites included in $\phi\in[0,\pi/4[$ produces a net positive result, i.e. a ferromagnetic moment that cancels with the contribution of the sites included in $\phi\in[\pi/4,\pi/2[$. The effect is very prominent for the $\{8,3\}$ lattice with $n_{l}=5$ layers [Fig. 5(c)], but we observed similar effects in other lattice geometries as well, also in the absence of zero modes [see Appendix D.1 for more examples].
It is tempting to identify our magnetic system as an altermagnet [64]. A corner stone of an altermagnet is compensated magnetism with spin-split bands. We can use the $C_{p}$ symmetry to identify a unit cell, $i$, with an orbital structure and corresponding Brillouin zone. Using this notation, the fermion creation operator can be written as $\hat{c}^{\dagger}_{i,n,\sigma}$ where $n$ denotes orbital $n$ in unit cell $i$. In this notation, the mean-field Hamiltonian reduces to:
$\hat{H}_{MF}=$ $\sum_{i,i^{\prime},n,n^{\prime},\sigma}\hat{c}^{\dagger}_{i,n,\sigma}T_{n,n}(i-i^{\prime})\hat{c}^{\phantom{\dagger}}_{i^{\prime},n^{\prime},\sigma}$
$+$ $\sum_{i,n,n^{\prime}\sigma}\sigma(-1)^{i}\hat{c}^{\dagger}_{i,n,\sigma}M_{n,n^{\prime}}\hat{c}^{\phantom{\dagger}}_{i,n^{\prime},\sigma}\,.$ (21)
In the above $M$ is a diagonal matrix that encodes the site magnetization in the unit cell. The factor $(-1)^{i}$ accounts for the global AFM. We can now carry out a Fourier transformation to obtain:
$\hat{H}_{MF}=\sum_{k\in\mathrm{MBZ},\sigma}$ $\left(\hat{\bm{c}}^{\dagger}_{k,\sigma},\hat{\bm{c}}^{\dagger}_{k+Q,\sigma}\right)$ (22)
$\begin{pmatrix}T(k)&\sigma M\\ \sigma M^{\dagger}&T(k+Q)\end{pmatrix}\begin{pmatrix}\hat{\bm{c}}^{\phantom{\dagger}}_{k,\sigma}\\ \hat{\bm{c}}^{\phantom{\dagger}}_{k+Q,\sigma}\end{pmatrix}\,.$
In the above $Q$ is the one-dimensional AFM wave vector, MBZ refers to the magnetic Brillouin zone and $\hat{\bm{c}}^{\dagger}_{k,\sigma}$ is a spinor accounting for the orbital degrees of freedom. Using the determinant identity based on the Schur complement, one will readily show that the energy spectrum does not depend on the spin index. As such, we cannot understand the observed global AFM in terms of an altermagnet.
Figure 7: (a) Average correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{b}}$ of the $n_{l}-1$ inner layers and (b) maximal correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{max}}$ as a function of the system size $\frac{1}{\sqrt{N_{s}}}$ at different strengths $\epsilon$ of the infinitesimal staggered field for the (I) $\{10,3\}$, (II) $\{8,3\}$ and (III) $\{8,8\}$ lattice.
Figure 8: Local correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to classical Néel state as obtained from spin wave approximation for the $\{8,3\}$ lattice with $n_{l}=4$ at $\epsilon=0.01$.
IV.3 Beyond the mean-field approximation
We now consider approximate as well as numerically exact methods that take into account fluctuations. We will first start with the spin wave theory of the Heisenberg model on hyperbolic lattices, and then use Monte Carlo methods for both the Hubbard and Heisenberg models.
Figure 9: (Ia)-(Ic) Finite-size scaling of the square of the AFM magnetization (order parameter), defined in Eq. (11), at $\beta t=100$ for different values of $U/t$ on the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice respectively (from left to right). On the $x$ axis we plot the square root of the inverse number of lattice sites $N_{s}$. The data points are extrapolated to $N_{s}\to\infty$ by a polynomial fit of second order and presented as a function of $U/t$ in the inset. We only present the extrapolated value for the $\{8,3\}$ lattice, since we have four data points to carry out the fit. We estimate that the systematic error is of the order of $0.005$ such that data points below this threshold should be interpreted with care. The data points for $U\to\infty$ refer to the Heisenberg model at $\beta J=100$, calculated by SSE. (IIa)-(IIc) Comparison of the squared magnetization, the squared magnetization of the bulk $\tilde{m}_{\textrm{AFM}}^{2}$ and the spin-spin correlation $\langle S_{\bm{r}}^{z}S_{0}^{z}\rangle$ of two sites with the largest distance to each other on the respective lattice, where $S_{0}^{z}$ is the site closest to $\phi=0$ (for the $\{10,3\}$-lattice this is negative since opposing sites lie on different sublattices). Here, we use $U/t=6$.
IV.3.1 Spin waves
The motivation to carry out this calculation stems from the fact that the majority of sites on the boundary of the hyperbolic lattice have a coordination number of two akin to a one-dimensional chain. We hence foresee that fluctuations on the edge will be greatly enhanced. To carry out the calculation we follow Refs. [65, 66], where spin-wave calculations were formulated for spin glasses. The method is summarized in Appendix B. To avoid zero modes, we include a small staggered field of magnitude $\epsilon$ and take the limit $\epsilon\rightarrow 0$.
In Fig. 7 we show the correction to the classical Néel state in the $\frac{1}{S}$-expansion as obtained from Eq. (28). Due to the absence of translation symmetry in the lattices, the corrections are site-dependent. The left panels show the average correction within the bulk $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{b}}$, where we defined the bulk as the $n_{l}-1$ inner layers, and the right panels show the maximal correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{max}}$, which can always be found on the sites closest to the angle $\phi=0$ in the outermost layer and all sites, that can be obtained by performing symmetry-allowed rotations.
As described in Appendix B the limit of $\epsilon\rightarrow 0$ and the thermodynamic limit have to be taken carefully. Taking the limit $\epsilon\rightarrow 0$ on finite lattices leads to divergencies in the correction [Fig. 7] and only in the thermodynamic limit $\frac{1}{\sqrt{N_{s}}}\rightarrow 0$ the correct results are recovered.
The average corrections in the bulk converge to values smaller than $\frac{1}{2}$ on the largest available lattice sizes for all three lattice geometries [Fig. 7(a)], meaning that the Néel state in the bulk is weakened, but still present in the case of spin $1/2$ degrees of freedom. The maximal correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{max}}$ on the other hand side is above $\frac{1}{2}$, such that it is large enough to destroy the magnetic moment on the corresponding sites.
In Fig. 8 the local correction of the $\{8,3\}$ lattice with $n_{l}=4$ layers at $\epsilon=0.01$ can be seen. The correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ depends strongly on the coordination number of the corresponding lattice site. In the bulk, where every site has $q=3$ neighbors, the correction is small. In the edge layer most sites are connected to only two neighboring sites, forming short one-dimensional sections. Therefore the correction on a given site is larger the greater the distance to the nearest site with $q=3$ neighbors. This result is consistent with the absence of long-range order in one dimension, as predicted by spin wave theory. Similar qualitative observations for the other lattice geometries can be found in Fig. 16 of Appendix D.2.
Figure 10: SSE calculated squared magnetization of the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice respectively (from left to right), at $\beta J=100$, together with the bulk and edge contribution. The data points were fitted with a second order polynomial.
Figure 11: Spin correlations $\langle S_{\bm{r}}^{z}S_{0}^{z}\rangle$ of the $\{10,3\}$ lattice for $n_{l}=3$ layers calculated by AFQMC at $U=1t$ and $\beta t=100$. As reference point $S_{0}^{z}$ we chose the site closest to $\phi=0$.
Figure 12: QMC data for the (Ia)-(Ic) uniform spin susceptibility $\chi_{S}$ and (IIa)-(IIc) charge susceptibility $\chi_{C}$ at $U=5t$ on the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice respectively (from left to right).
IV.3.2 Quantum Monte Carlo
The aim of this section is to use numerically exact approaches, AFQMC and SSE, to check the validity of the above statements.
Figures 9(Ia)-9(Ic) plot the staggered spin-spin-correlations as defined in Eq. (11) as a function of $\frac{1}{\sqrt{N_{s}}}$. Unless mentioned otherwise, we consider a finite, but low temperature scale $\beta t=100$ and fit the data to the form $m^{2}_{N_{s}\rightarrow\infty}+aN_{s}^{-\frac{1}{2}}+bN_{s}^{-1}$. It is notoriously hard to extrapolate the staggered magnetization to the thermodynamic limit. First, since the number of sites grow rapidly with $n_{l}$ we only have a limited set of data. Second, we measure the square of the order parameter such that when the staggered magnetization is small (i.e. at weak coupling or in the proximity of a quantum phase transition) very large lattices are required to obtain reliable results. We note that a pinning field approach [48] that was used for the Hubbard model on the honeycomb lattice may be a alternative and promising method for future investigations.
For the $\{10,3\}$ and $\{8,3\}$ lattices [Figs. 9(Ia) and 9(Ib)] our data at $\beta t=100$ ($\beta J=100$) is representative of the ground-state properties for the considered values of $n_{l}$. For the $\{8,3\}$ lattice ordering sets in at $U_{c}\simeq 2t$. Upon inspection we see that the $\{10,3\}$ lattice is still disordered at $U/t=3$. Hence, the trends in the value of $U_{c}$, $(U_{c}^{\{10,3\}}>U_{c}^{\{8,3\}}$) between the two lattices match our mean-field analysis. It is beyond the scope of this article to determine critical exponent of the order parameter: $m\simeq(U-U_{c})^{\beta}$. In Figs. 9(Ia) and 9(Ib), we equally consider the Heisenberg model. For this model, the SSE allows us to reach much larger lattices such that the extrapolation is more robust.
In Fig. 9(Ic) we consider the flat band system $\{8,8\}$. With the fermion QMC we can only access three system sizes for the extrapolation. Hence, we do not attempt to determine $m^{2}_{N_{s}\rightarrow\infty}$ for this lattice. Furthermore, temperature effects are much larger for this flat band system. Nevertheless, the data is consistent with ordering at large values of $U/t$ and for the Heisenberg model.
In Figs. 9(IIa)-9(IIc), we also present the AFM magnetization of the bulk $\tilde{m}^{2}_{AFM}$ by only summing over sites of the $n_{l}-1$ inner layers. The bulk magnetization is larger than the total magnetization including the edge, consistent with our results from spin wave theory. This becomes more obvious when we consider the bulk and edge contribution to the magnetization calculated by SSE, as presented in Fig. 10. This method allows a more accurate study of the magnetization in the strong-coupling limit as larger lattices can be simulated. While the edge magnetization is slightly lower than the total magnetization, the bulk contribution is larger by a factor of approximately $\sqrt{2}$ for all considered lattices. This specifies our findings from spin wave theory that fluctuations are enhanced on the boundary. The SSE calculations suggest more precisely, that there is still long range order on the edge. Finally, we note that due to the sub-extensiveness of the bulk, the magnetization is determined by the edge contribution. This is supported by the SSE data of Fig. 10 where the total and edge data scale to the same value in the large $n_{l}$ limit.
The global AFM is a weak-coupling effect, such that we use the Hubbard model and AFQMC for investigations. To compare the results from mean-field with QMC, we present spin correlations plotted in the Poincaré disk of the $\{10,3\}$ lattice as shown in Fig. 11. Again considering the weak-coupling regime, similar patterns of alternating positive and negative correlations can be observed in the third layer. This can be directly compared to the mean-field results of the order parameter for the $\{10,3\}$ lattice in Fig. 13(a), see Appendix. For the presentation of the QMC results, we choose the $\{10,3\}$ lattice, since it offers the clearest visualization of the observed effect. Note that the considered value of $U/t$ is smaller than the estimated $U_{c}$. Nevertheless, the observed spin fluctuations support the notion that in the symmetry-broken phase at large $U/t$ the $C_{10}$ and time-reversal symmetry are broken, but $C_{10}\mathcal{T}$ remains a symmetry of the state.
In order to support the mean-field observation that the system is in an AFM Mott insulating state for strong electronic correlations, we present the spin and charge susceptibility for $U=5t$ on the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice in Figs. 12(Ia)-12(Ic) and 12(IIa)-12(IIc) respectively. As expected for an AFM state spin waves lead to a spin susceptibility $\chi_{S}$ that converges to a finite value as $T\to 0$. The data shows that there is a buildup of low energy states as one increases $n_{l}$. On the other hand, the charge susceptibility $\chi_{C}$ decays exponentially, as expected for a charged, gaped system. Figures 12(IIa)-12(IIc) show that as $n_{l}$ grows the charge susceptibility remains activated in the low-temperature limit.
V Discussion and Conclusion
The physics of the half-filled Hubbard and Heisenberg models on bipartite lattices has been investigated in great details for Bravais lattices. Owing to a theorem by Lieb [51], the ground state on a finite lattice is a spin singlet. The lowest energy state in each total spin sector builds the Anderson tower of states [67], that collapses in the thermodynamic limit. This is the mechanism that allows for the broken-symmetry quantum AFM ground state without violating the aforementioned Lieb theorem in dimensions greater or equal to two. In one dimension, the Heisenberg and Hubbard models, are critical and are described by an SO(4) Wess-Zumino-Witten theory.
The notion of thermodynamic limit for the hyperbolic lattices differs from the Euclidean case: for the hyperbolic lattices the boundary dominates over the bulk. As a consequence, the very notion of dimensionality and role of fluctuations can be questioned. While our mean-field calculations for the Hubbard model show long-range order on the boundary and in the bulk, a spin wave calculation for the Heisenberg model shows that fluctuations inhibit order at the boundary. As demonstrated by our QMC simulations for both the strong-coupling Hubbard and Heisenberg models, this turns out to be an artifact of the spin wave approximation. It is known that spin waves fail to predict that weakly coupling chains with non-frustrating interactions is a relevant perturbation [68] that leads to long range magnetic order. However, even if the ground state is ordered, it would be of interest to investigate the spin dynamical response on the boundary, with the aim to assess whether proximity to one-dimensional physics is visible at high frequencies as in Ref. [68].
Another direction of research is to consider frustration. It is known that weak frustrating couplings between chains do not necessarily lead to magnetic ordering [69], thereby offering the possibility of realizing different phases in the bulk and on the boundary of hyperbolic lattices.
As for the honeycomb lattice, the Dirac-like character of the DOS close to the Fermi level of the $\{10,3\}$ and $\{8,3\}$ lattices lead to a finite critical interaction strength $U_{c}$ for the onset of AFM ordering. At the mean-field level the order parameter exponent takes the value $\beta=1$ as for the Euclidean case. On Euclidean lattices, great progress has been achieved in the understanding of the Gross-Neveu transition, by using renormalization group invariant quantities and finite size scaling to analyze the data [47, 48, 49]. For the hyperbolic lattices our analysis is very rudimentary due to the very difficulty of defining a continuum limit and the accompanying critical theory. We find this to be an interesting line of future research. Nevertheless, our data is consistent with a finite-$U$ Mott transition.
The thermodynamic properties of the AFM Mott insulating state are the very same as in the Euclidean case. At low temperatures and in the large $n_{l}$ limit, the charge susceptibility is activated and the spin susceptibility constant. We note that the notion of Goldstone modes is a property of the symmetry group and not of the lattice structure. As such it is not surprising to observe a finite spin susceptibility in the low-temperature limit.
In the weak-coupling limit patterns of global AFM occur and can be observed in both mean-field and QMC results. Following the probability distribution of the low-lying energy eigenstates, macroscopic ferromagnetic moments are formed, that are compensated in sum. These patterns all have in common, that they break the $C_{p}$ (down to $C_{p/2}$) and time-reversal symmetry of the Hamiltonian. Additionally the magnetic patterns are invariant under a combination of the original rotation group of the lattice and time-reversal symmetry $C_{p}\mathcal{T}$. Global AFM occurs e.g. also in finite graphene samples with zig-zag edges [70, 71, 72, 73, 74, 75, 76] or in more complex pattern, when the sample is subject to a strain [77, 63, 78].
The absence of negative sign problem in QMC simulations hinges on symmetries (e.g. combined time-reversal and a U(1) conservation) [79, 80] that are not broken by the hyperbolic geometry. Hence all models that are amenable to negative sign-free QMC simulations such as SU(N) Hubbard-Heisenberg models [81], SU(N) Kondo lattices [82, 61], and Su-Schrieffer-Heeger Hamiltonians [83, 84] can be studied. It remains to be seen if novel phenomena can be observed due the hyperbolic geometry.
Acknowledgements.
The authors thank Igor Boettcher and Ronny Thomale for discussions. Special thanks to Igor Boettcher for providing a Mathematica notebook that generates adjacency matrices for hyperbolic lattices. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at the Leibniz Supercomputing Centre (www.lrz.de). FA and JI thank the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490), GR and AG the DFG funded SFB 1170 on Topological and Correlated Electronics at Surfaces and Interfaces. FA acknowledges financial support from the German Research Foundation (DFG) under the grant AS 120/16-1 (Project number 493886309) that is part of the collaborative research project SFB Q-M&S funded by the Austrian Science Fund (FWF) F 86.
Appendix A Lattice sizes
Table 1 summarizes the lattices used in the main text and their respective number of lattice sites $N_{s}$ as a function of the number of layers $n_{l}$.
$\{10,3\}$$\{8,3\}$$\{8,8\}$
$n_{l}=1$ 10 8 8
$2$ 80 48 56
$3$ 490 200 392
$4$ 2880 768 2744
$5$ 16810 2888
$6$ 10800
Table 1: Number of lattice sites for considered lattice geometries as a function of the number of layers $n_{l}$.
Appendix B Spin wave approximation
[…]
$\hat{H}=S\sum_{\bm{i}}z_{\bm{i}}\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}+\frac{S}{2}\sum_{\bm{i},\bm{j}}T_{\bm{i},\bm{j}}(\hat{b}_{\bm{i}}^{\phantom{\dagger}}\hat{b}_{\bm{j}}^{\phantom{\dagger}}+\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{j}}^{\dagger})-S^{2}N_{b}\,.$ (23)
Here we defined the site-dependent coordination number $z_{\bm{i}}=\sum_{\bm{j}}T_{\bm{i},\bm{j}}$. We can write the equations of motions for the boson operators in a matrix representation as
$\frac{d}{dt}\hat{\bm{b}}=i\begin{pmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{pmatrix}\hat{\bm{b}}$ (24)
with
$\hat{\bm{b}}^{\dagger}$ $=$ $(\hat{b}_{\bm{i}=1}^{\dagger},\dots,\hat{b}_{\bm{i}=N}^{\dagger},\hat{b}_{\bm{i}=1}^{\phantom{\dagger}},\dots,\hat{b}_{\bm{i}=N}^{\phantom{\dagger}})\,,$
$P_{\bm{i},\bm{j}}$ $=$ $Sz_{\bm{i}}\delta_{\bm{i},\bm{j}}\,,\quad Q_{\bm{i},\bm{j}}=ST_{\bm{i},\bm{j}}\,.$ (25)
In order to diagonalize the Hamiltonian of Eq. (23) we perform a Bogoliubov transformation
$\begin{pmatrix}-\Omega&0\\ 0&\Omega\\ \end{pmatrix}=\begin{pmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{pmatrix}^{-1}\begin{pmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{pmatrix}\begin{pmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{pmatrix}\,.$ (26)
To ensure, that the new boson operators $\hat{d}_{\bm{i}}^{\dagger}$, that we obtain from $\hat{b}_{\bm{i}}^{\dagger}=\bar{f}_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\phantom{\dagger}}+g_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\dagger}$, still fulfill the canonical commutation relations, we need to impose a condition on the entries of $f$ and $g$
$\begin{pmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{pmatrix}\begin{pmatrix}-1&0\\ 0&1\\ \end{pmatrix}\begin{pmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{pmatrix}^{\dagger}=\begin{pmatrix}-1&0\\ 0&1\\ \end{pmatrix}\,.$ (27)
We now define a scalar product $\langle\bm{x},\bm{y}\rangle_{B}=\bm{x}^{\dagger}\left(\begin{smallmatrix}-1&0\\ 0&1\\ \end{smallmatrix}\right)\bm{y}$. The matrix $M=\left(\begin{smallmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{smallmatrix}\right)$ is hermitian with respect to this scalar product $\langle M\bm{x},\bm{y}\rangle_{B}=\langle\bm{x},M\bm{y}\rangle_{B}$ and the Bogoliubov transformation $U=\left(\begin{smallmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{smallmatrix}\right)$ is unitary $\langle U\bm{x},U\bm{y}\rangle_{B}=\langle\bm{x},\bm{y}\rangle_{B}$. After diagonalizing the matrix $M$, the condition on the Bogoliubov transformation of Eq. (27) can be satisfied by ensuring, that all eigenvectors of $M$ are orthonormal with respect to the scalar product $\langle\cdot,\cdot\rangle_{B}$. If the norm of an eigenvector $\langle\bm{x}_{n},\bm{x}_{n}\rangle_{B}>0$ is nonzero, it can also be shown, that the corresponding eigenvalue is real. In order to avoid eigenvectors with a vanishing $B$ norm we add an infinitesimal staggered magnetic field $\hat{H}_{\epsilon}=-\epsilon\sum_{\bm{i}}(-1)^{\bm{i}}\hat{S}_{\bm{i}}^{z}$. After the calculation, limits have to be taken carefully, by first taking the thermodynamic limit and then the limit of $\epsilon\rightarrow 0$. The site-dependent correction to the classical AFM state can be calculated by the mean boson occupation number at $T=0$
$\langle\hat{S}_{\bm{i}}^{z}\rangle=S-\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle=S-\sum_{\bm{j}}|f_{\bm{i},\bm{j}}|^{2}\,.$ (28)
Appendix C Spin susceptibility
In this appendix we show that the assumption of partial particle-hole symmetry and a finite DOS at the Fermi surface, suffices to argue a logarithmic divergence of the staggered spin susceptibility. This calculation does not depend on the lattice symmetry and would also hold for disordered systems where the disorder does not break the said symmetry. The tight-binding part of the Hamiltonian is invariant under a partial particle-hole transformation defined as
$\hat{P}^{-1}_{\sigma}\hat{c}_{\bm{i},\sigma^{\prime}}^{\dagger}\hat{P}_{\sigma}=\delta_{\sigma,\sigma^{\prime}}(-1)^{\bm{i}}\hat{\gamma}_{\bm{i},\sigma^{\prime}}^{\phantom{\dagger}}+\left(1-\delta_{\sigma,\sigma^{\prime}}\right)\hat{\gamma}_{\bm{i},\sigma^{\prime}}^{\dagger}\,.$ (29)
Without loss of generality we can consider the staggered spin susceptibility in $x$-direction
[…]
As apparent, the partial particle-hole transformation maps the staggered susceptibility in the particle-hole channel to the uniform one in the particle-particle channel. The latter is nothing but the Cooper instability, that we now show to be present on any graph with finite DOS at the Fermi energy. To do so, we introduce a canonical transformation $U$, that diagonalizes the Hamiltonian
$\hat{H}_{t}=\sum_{\bm{i},\bm{j},\sigma}h_{\bm{i},\bm{j}}\hat{\gamma}_{\bm{i},\sigma}^{\dagger}\hat{\gamma}_{\bm{j},\sigma}^{\phantom{\dagger}}=\sum_{x}\lambda_{x}\hat{\eta}_{x,\sigma}^{\dagger}\hat{\eta}_{x,\sigma}^{\phantom{\dagger}}$ (31)
with new fermionic operators $\hat{\eta}_{x,\sigma}^{\dagger}=\sum_{\bm{i}}\hat{\gamma}_{\bm{i},\sigma}^{\dagger}U_{\bm{i},x}$. Applying this transformation yields
[…]
×$\times\sum_{x,y}f_{x,y}e^{\tau(\lambda_{x}+\lambda_{y})}\langle\hat{\eta}_{x,\uparrow}^{\dagger}\hat{\eta}_{x,\uparrow}^{\phantom{\dagger}}\rangle\langle\hat{\eta}_{y,\downarrow}^{\dagger}\hat{\eta}_{y,\downarrow}^{\phantom{\dagger}}\rangle$
$=$ $\frac{1}{N_{s}}\sum_{x,y}\frac{f_{x,y}}{\lambda_{x}+\lambda_{y}}\left[\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}+\tanh{\left(\frac{\beta}{2}\lambda_{y}\right)}\right]$
with $f_{x,y}=\sum_{\bm{i},\bm{j}}(U^{\dagger})_{x,\bm{i}}U_{\bm{j},x}(U^{\dagger})_{y,\bm{i}}U_{\bm{j},y}$. If the DOS is finite at zero energy, the Fermi level here, we can show, that the term with $x=y$ has a logarithmic divergency
$\chi^{x}(Q)$ $=$ $\frac{1}{N_{s}}\sum_{x}\frac{f_{x,x}}{\lambda_{x}}\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}+R$ (33)
[…]
where all other terms in the sum with $x\neq y$ are summarized in the term $R$. The last approximation is in the low-temperature limit and $W$ is a high-energy cutoff. In principle terms could appear in $R$, that lead to the cancellation of the logarithmic divergency under certain conditions. But in general the occurrence of an instability towards AFM ordering is generic even in the absence of translational symmetry.
Appendix D Additional data
In the following section we present some additional data.
D.1 Mean-field approximation
Figure 13: (I) Local DOS $\rho(E=0,\bm{i})$ at zero energy in the non-interacting limit $U=0$ at $\delta=0.01$ and (II) magnetization $m_{\bm{i}}^{z}$ as obtained from mean-field approximation for the $\{10,3\}$ lattice and interaction strength $U=0.8t$.
Figure 14: Same as in Fig. 13, but for the $\{8,3\}$ lattice and interaction strength (IIa) $U=0.4t$ and (IIb) $U=0.58t$.
In Sec. IV.2 we showed the spatial distribution of the magnetization $m_{\bm{i}}^{z}$ in the weak-coupling limit for the $\{8,3\}$ lattice with $n_{l}=5$ layers. In the upper panels of Figs. 13-15 we plot the local DOS at zero energy
$\rho(E,\bm{i})=-\frac{1}{\pi N_{s}}\sum_{n}|U_{\bm{i},n}|^{2}\mathrm{Im}G^{R}(n,E)$ (34)
for the different lattice geometries both with $n_{l}=3$ and $n_{l}=4$ layers.
Figure 15: Same as in Fig. 13, but for the $\{8,8\}$ lattice and interaction strength $U=0.1t$.
Here we defined a canonical transformation $U$, that diagonalizes the mean-field Hamiltonian (4). They all show a unique pattern of low and high DOS, that is compatible with the $C_{p}$ symmetry of the lattices. The lower panels of Figs. 13-15 show the corresponding local order parameter in real-space, where again local ferromagnetic and AFM moments emerge in the different sectors of the lattice. But in sum of all lattice sites only the AFM moment remains.
D.2 Spin wave approximation
The results for the site-dependent correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to the classical Néel state obtained from the spin wave approximation for the $\{10,3\}$ and $\{8,3\}$ lattice in Figs. 16(I) and 16(II) are similar to the results for the $\{8,3\}$ lattice described in the main text [Fig. 8]. For the $\{8,8\}$ lattice the results differ slightly due to the lattice geometry. The correction on many sites in the outermost layer is smaller than for the other two lattices, since the distance from one site on the edge to another site on the edge is in general smaller. This is a consequence of the high coordination number $q=8$, since sites in the bulk can have direct connections to the edge. The sites on the edge also form short one-dimensional chains, but larger effective interactions are present in comparison to the other two considered lattice geometries.
Figure 16: Local correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to classical Néel state as obtained from spin wave approximation for the (I) $\{10,3\}$, (II) $\{8,3\}$ and (III) $\{8,8\}$ lattice and varying system size $n_{l}$ at $\epsilon=0.01$.
D.3 Quantum Monte Carlo
Figure 17: AFM order parameter for the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice obtained by QMC calculations at $\beta t=100$ (filled symbols) and $\beta t=25$ (unfilled symbols).
In order to supplement the picture given in Figs. 9 and 11, we present the order parameter $m_{AFM}^{z}$ as a function of $U$ in Fig. 17 for two different temperatures $\beta t=25$ and $\beta t=100$. We see that the $\{8,8\}$ lattice exhibits stronger temperature effects than the other lattices. This lattice shows a strong sensitivity to temperature already in the weak coupling regime, as can be seen in Fig. 17(c). As discussed before, this effect is also present in the results from a mean-field analysis [Fig. 4(a)].

### v2

@@
-Hubbard and Heisenberg models on hyperbolic lattices - Metal-insulator transitions, global antiferromagnetism and enhanced boundary fluctuations
+Hubbard and Heisenberg models on hyperbolic lattices: Metal-insulator transitions, global antiferromagnetism, and enhanced boundary fluctuations
 Abstract
-We study the Hubbard and Heisenberg models on hyperbolic lattices with open boundary conditions by means of mean-field approximations, spin-wave theory and quantum Monte Carlo (QMC) simulations.
-For the Hubbard model we use the auxiliary-field approach and for Heisenberg systems the stochastic series expansion algorithm.
-We concentrate on bipartite lattices where the QMC simulations are free of the negative sign problem.
-The considered lattices are characterized by a Dirac like density of states, Schläfli indices $\{p,q\}=\{10,3\}$ and $\{8,3\}$, as well as by flat bands, $\{8,8\}$.
-The Dirac density of states cuts off the logarithmic divergence of the staggered spin susceptibility and allows for a finite $U$ semi-metal to insulator transition.
-This transition has the same mean-field exponents as for the Euclidean counterpart.
+We study the Hubbard and Heisenberg models on hyperbolic lattices with open boundary conditions by means of mean-field approximations, spin-wave theory, and quantum Monte Carlo (QMC) simulations.
+For the Hubbard model we use the auxiliary-field approach and for Heisenberg systems the stochastic series expansion algorithm and concentrate on bipartite lattices where the QMC simulations are free of the negative sign problem.
+The hyperbolic lattices have an extensive number of sites on the boundary, such that one has to distinguish between bulk and total density of states.
+The considered lattices are characterized by a Dirac-like total density of states, Schläfli indices $\{p,q\}=\{10,3\}$ and $\{8,3\}$, as well as by flat bands, $\{8,8\}$.
+The Dirac total density of states cuts off the logarithmic divergence of the staggered spin susceptibility and allows for a finite $U$ metal-to-insulator transition.
+This transition has the same mean-field exponents as for the Gross-Neveu transition in Euclidean space.
+We argue that this transition is induced by the open-boundary conditions and that it will be absent in the periodic case.
 In the presence of flat bands we observe the onset of magnetic ordering at any finite $U$.
+This conclusion holds even though the bulk density of states is constant at the Fermi energy for the three considered lattices.
 The magnetic state at intermediate coupling can be described as a global antiferromagnet.
-It breaks the $C_{p}$ rotational and time reversal symmetries but remains invariant under combined $C_{p}\mathcal{T}$ transformations.
+It breaks the $C_{p}$ rotational and time-reversal symmetries but remains invariant under combined $C_{p}\mathcal{T}$ transformations.
 The state is characterized by macroscopic ferromagnetic moments, that globally cancel.
-We observe that fluctuations on the boundary of the system are greatly enhanced: while spin wave calculations predict the breakdown of antiferromagnetism on the boundary but not in the bulk, QMC simulations show a marked reduction of the staggered moment on the edge of the system.
+We observe that fluctuations on the boundary of the system are greatly enhanced: While spin-wave calculations predict the breakdown of antiferromagnetism on the boundary but not in the bulk, QMC simulations show a marked reduction of the staggered moment on the edge of the system.
 I Introduction
@@
 Also in condensed matter physics the popularity of hyperbolic lattices is increasing.
-The topics of interest are both single-particle physics, like the electronic band structure [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] – also under the influence of a magnetic field and topological phenomena – and correlation effects [28, 29, 26].
+The topics of interest are both single-particle physics, like the electronic band structure [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]—also under the influence of a magnetic field and topological phenomena—and correlation effects [28, 29, 26].
 The motivation can partially be attributed to recent experimental realizations of hyperbolic lattices in circuit quantum electrodynamics and electric circuits [30, 31].
-In this paper we study the Hubbard and Heisenberg models on three exemplary regular tessellations of the hyperbolic plane, that exhibit different electronic properties in the non-interacting limit.
-We use a Hartree-Fock approximation, a spin wave analysis, auxiliary-field and stochastic series expansion quantum Monte Carlo (QMC) methods to investigate the models as a function of the interaction strength and system size.
+In this paper we study the Hubbard and Heisenberg models on three exemplary regular tessellations of the hyperbolic plane, that exhibit different electronic properties in the noninteracting limit.
+We use a Hartree-Fock approximation, a spin-wave analysis, auxiliary-field and stochastic series expansion quantum Monte Carlo (QMC) methods to investigate the models as a function of the interaction strength and system size.
 Since all considered lattices are bipartite, we can perform QMC simulations without a negative sign problem for both the Hubbard and Heisenberg models.
-The electronic properties of the three lattices resemble a Dirac system with a linear density of states (DOS) around the Fermi level ($\{10,3\}$ and $\{8,3\}$ lattice), and a flat-band system with a high DOS at the Fermi level ($\{8,8\}$ lattice).
+There is a plethora of exotic phases of matter, such as spin liquids, that have been realized on two-dimensional bipartite lattices [32, 33, 34, 35].
+This provides a further motivation to study correlation effects on bipartite hyperbolic lattices.
+We will concentrate on systems with open boundary conditions.
+This is important since the number of boundary sites is extensive, such that bulk and total density of states (DOS) differ.
 The mean-field approximation of the Hubbard model on the $\{10,3\}$ lattice is already studied in Ref.
-[26], and we complement these results with both spin wave calculations and first-time fermionic QMC studies on the hyperbolic lattice.
-We will compare the spatial distribution of the magnetic ordering in the weak-coupling and strong-coupling limits, the transition from a semi-metal to insulator with antiferromagnetic (AFM) ordering as a function of the interaction strength and thermodynamic properties.
+[26], and we complement these results with both spin-wave calculations and first-time fermionic QMC studies on the hyperbolic lattice.
 The remainder of the paper is organized as follows.
@@
 V.
-In the Appendix additional numerical data can be found.
+Additional numerical data as well as details and derivations of the methods used in this paper can be found in the Appendix.
 II Hyperbolic Lattices, Hubbard and Heisenberg Models
@@
 Hyperbolic lattices with a constant negative curvature fulfill the condition $(p-2)(q-2)>4$.
-In contrast, lattices with $(p-2)(q-2)=4$ lie within the flat Euclidean plane and lattices with $(p-2)(q-2)<4$ in a positively curved plane.
-For the Euclidean plane exactly three solutions exist – the square lattice $\{4,4\}$, the honeycomb lattice $\{6,3\}$ and the triangular lattice $\{3,6\}$, whereas for the hyperbolic case infinitely many pairs $\{p,q\}$ exist that fulfill the condition [33].
-We concentrate on finite size hyperbolic lattices with open boundary conditions.
+In contrast, lattices with $(p-2)(q-2)=4$ lie within the flat Euclidean plane and lattices with $(p-2)(q-2)<4$ on a sphere, that has a constant positive curvature.
+For the Euclidean plane exactly three solutions exist—the square lattice $\{4,4\}$, the honeycomb lattice $\{6,3\}$ and the triangular lattice $\{3,6\}$, and for the sphere five—the five Platonic solids, whereas for the hyperbolic case infinitely many pairs $\{p,q\}$ exist that fulfill the condition [37].
+We concentrate on finite-size hyperbolic lattices with open boundary conditions.
 The lattice size can be increased by starting with a single $p$-gon and successively adding shells or layers of $p$-gons.
@@
 1 the $\{8,3\}$ lattice is shown with up to $n_{l}=3$ layers.
-In order to properly depict the lattices, the Poincaré disk model is used.
+To properly depict the lattices, the Poincaré disk model is used.
 All lattice sites are projected into the open unit disk and all edges, connecting two nearest-neighboring sites, are geodesic lines.
-Those geodesics are circular arcs in the Poincaré disk model, that would hit the unit circle perpendicular and they all have the same length within the hyperbolic metric  [^1] The metric in the hyperbolic plane with a constant negative curvature of -1 is given by $ds^{2}=4\frac{dx^{2}+dy^{2}}{(1-x^{2}-y^{2})^{2}}$.
+Those geodesics are circular arcs in the Poincaré disk model, that would hit the unit circle perpendicularly and they all have the same length within the hyperbolic metric  [^1] The metric in the hyperbolic plane with a constant negative curvature of $-1$ is given by $ds^{2}=4\frac{dx^{2}+dy^{2}}{(1-x^{2}-y^{2})^{2}}$.
 Here, $x$ and $y$ are the coordinates within the open unit circle in the Poincaré disk model.
@@
 For open boundary conditions the ratio of lattice sites at the edge to the total number of lattice sites converges to a finite value, so the behavior of the edge sites is not negligible even for large system sizes [3, 6, 24, 27].
-If the bulk properties are of interest, the lattice can be successively increased and the calculation can only be done for the inner bulk part, neglecting the outer shells [3, 4].
+If the bulk properties are of interest, then the lattice can be successively increased and the calculation can only be done for the inner bulk part, neglecting the outer shells [3, 4].
 Another possibility is to use periodic boundary conditions.
 Some classes of hyperbolic tilings can be placed on compact surfaces with higher genus $g\geq 2$ with periodic boundary conditions [41, 8, 28, 18, 19].
-But in this work we use open boundary conditions and also take into account the boundary contributions.
-Fig.
-1 also shows, that the $\{8,3\}$ lattice is bipartite with sublattice $A$ and $B$, like all lattices in this work.
+In this work we use open boundary conditions and also take into account the boundary contributions.
+Table 1: Number of lattice sites for considered lattice geometries as a function of the number of layers $n_{l}$.
+$\{10,3\}$$\{8,3\}$$\{8,8\}$
+$n_{l}=1$ 10 8 8
+$2$ 80 48 56
+$3$ 490 200 392
+$4$ 2880 768 2744
+$5$ 16810 2888
+$6$ 10800
+Figure 1 also shows, that the $\{8,3\}$ lattice is bipartite with sublattice $A$ and $B$, like all lattices in this work.
 In general for even $p$ the resulting lattices are bipartite [27].
-In Table 1 of the Appendix we provide a list of the considered lattice sizes and their respective number of lattice sites $N_{s}$.
+In Table 1 we provide a list of the considered lattice sizes and their respective number of lattice sites $N_{s}$.
 II.2 Hubbard and Heisenberg models
@@
 $\hat{H}$ $=$ $\hat{H}_{t}+\hat{H}_{U}\,,$
-$\hat{H}_{t}$ $=$ $-t\sum_{\langle\bm{i},\bm{j}\rangle}\sum_{\sigma=1}^{N_{\sigma}}\left(\hat{c}_{\bm{i},\sigma}^{\dagger}\hat{c}_{\bm{j},\sigma}^{\phantom{\dagger}}+\mathrm{h.c.}\right)\,,$
-$\hat{H}_{U}$ $=$ $\frac{U}{N_{\sigma}}\sum_{\bm{i}}\left(\sum_{\sigma=1}^{N_{\sigma}}\hat{n}_{\bm{i},\sigma}-\frac{N_{\sigma}}{2}\right)^{2}\,.$ (1)
+$\hat{H}_{t}$ $=$ $-t\sum_{\langle\bm{i},\bm{j}\rangle}\sum_{\sigma=1}^{N}\left(\hat{c}_{\bm{i},\sigma}^{\dagger}\hat{c}_{\bm{j},\sigma}^{\phantom{\dagger}}+\mathrm{H.c.}\right)\,,$
+$\hat{H}_{U}$ $=$ $\frac{U}{N}\sum_{\bm{i}}\left(\sum_{\sigma=1}^{N}\hat{n}_{\bm{i},\sigma}-\frac{N}{2}\right)^{2}\,.$ (1)
 The first term describes the hopping of fermions on bonds $b=\langle\bm{i},\bm{j}\rangle$ connecting nearest-neighboring sites $\bm{i}$ and $\bm{j}$ with hopping strength $t$.
 The operator $\hat{c}_{\bm{i},\sigma}^{\dagger}$ creates a fermion in a Wannier state centered at site $\bm{i}$ and with $z$ component of spin equal to $\sigma$.
-We will consider $N_{\sigma}=2$ (spin $S=\frac{1}{2}$) flavored fermions.
+We will consider $N=2$ (spin $S=\frac{1}{2}$) flavored fermions.
 The fermions interact via a repulsive on-site Hubbard potential with strength $U$, as described in the second term.
@@
 This effect can be observed for the available regular bipartite graphs in the Euclidean plane, the square lattice $\{4,4\}$ and the honeycomb lattice $\{6,3\}$.
-On both lattices the repulsive Hubbard model at half-filling undergoes a transition from a (semi-)metal to an AFM insulator in the limit of strong on-site interactions.
+On both lattices the repulsive Hubbard model at half-filling undergoes a transition from a (semi-)metal to an antiferromagnetic (AFM) insulator in the limit of strong on-site interactions.
 The Fermi surface of the half-filled square lattice exhibits van Hove singularities leading to a logarithmically divergent DOS at the Fermi surface.
@@
 Numerical studies support the presence of long-range AFM order for any finite interaction $U$ [45, 47, 48, 49].
-On the honeycomb lattice the band structure of the non-interacting case exhibits two Dirac cones and the Fermi surface is concentrated on two points with a vanishing DOS.
+On the honeycomb lattice the band structure of the noninteracting case exhibits two Dirac cones and the Fermi surface is concentrated on two points with a vanishing DOS.
 A finite critical on-site repulsion $U_{c}$ is needed for the formation of long-range AFM ordering [50].
-Numerical studies locate a continuous and direct semi-metal to insulator transition at $U_{c}/t\approx 3.8$ [47, 48, 49].
+Numerical studies locate a continuous and direct semimetal to insulator transition at $U_{c}/t\approx 3.8$ [51, 52, 53].
 This transition belongs to the Gross-Neveu universality class [54, 52, 53].
@@
 $S_{\mathrm{tot}}$ is the total spin and the theorem was proven for any bipartite collection of sites, meaning no explicit periodicity or dimensionality is assumed in the proof of the theorem.
-Furthermore in the strong-coupling limit the model maps onto the Heisenberg model with an AFM coupling [52], since the mapping is independent of the lattice.
+Furthermore, in the strong-coupling limit the model maps onto the Heisenberg model with an AFM coupling [56], since the mapping is independent of the lattice.
 The specific form of the model reads:
-$\hat{H}=J\sum_{\langle\bm{i},\bm{j}\rangle}\hat{\bm{S}}_{\bm{i}}\cdot\hat{\bm{S}}_{\bm{j}}$ (3)
+$\hat{H}=J\sum_{\langle\bm{i},\bm{j}\rangle}\hat{\bm{S}}_{\bm{i}}\cdot\hat{\bm{S}}_{\bm{j}}\,,$ (3)
 with $J=\frac{4t^{2}}{U}$ and $\hat{\bm{S}}_{\bm{i}}=\frac{1}{2}\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{\bm{i},\sigma}\bm{\sigma}_{\sigma,\sigma^{\prime}}\hat{c}^{\phantom{\dagger}}_{\bm{i},\sigma^{\prime}}$
@@
 where $\langle\dots\rangle_{\phi}=\frac{\mathop{\mathrm{Tr}}[\dots e^{-\beta\hat{H}_{\mathrm{MF}}}]}{\mathop{\mathrm{Tr}}[e^{-\beta\hat{H}_{\mathrm{MF}}}]}$ is the expectation value with respect to $\hat{H}_{\mathrm{MF}}$ and $\beta=T^{-1}$ the inverse temperature.
-The fields $\phi_{\bm{i}}$ are equivalent to the local magnetic order parameter and due to the absence of translation symmetry in the considered lattices, we allow the fields to vary from site to site.
+The fields $\phi_{\bm{i}}$ are equivalent to the local magnetic order parameter and due to the absence of translation symmetry in the considered lattices with open boundary conditions, we allow the fields to vary from site to site.
 In the following, Eq.
 (5) will be solved self-consistently.
-In Appendix C we show that, due to particle-hole symmetry, the staggered spin susceptibility diverges logarithmically in the low-temperature limit, the prefactor being the DOS at the Fermi energy.
+In Appendix B we show that, due to particle-hole symmetry, the staggered spin susceptibility diverges logarithmically in the low-temperature limit, the prefactor being the DOS at the Fermi energy.
 Hence, we foresee AFM ordering and define the staggered order parameter
 $m^{z}=\frac{1}{N_{s}}\sum_{\bm{i}}(-1)^{\bm{i}}m^{z}_{\bm{i}}\,.$ (6)
-III.2 Spin wave approximation
-We continue with a spin wave approximation. In the strong-coupling limit the Hubbard model maps onto the AFM Heisenberg model independently of the lattice geometry [52]
+III.2 Spin-wave approximation
+We continue with a spin-wave approximation. In the strong-coupling limit the Hubbard model maps onto the AFM Heisenberg model independently of the lattice geometry [56]
 $\hat{H}=\sum_{\langle\bm{i},\bm{j}\rangle}\hat{\bm{S}}_{\bm{i}}\cdot\hat{\bm{S}}_{\bm{j}}=\frac{1}{2}\sum_{\bm{i},\bm{j}}T_{\bm{i},\bm{j}}\hat{\bm{S}}_{\bm{i}}\cdot\hat{\bm{S}}_{\bm{j}}\,.$ (7)
@@
 Since we start from an AFM, we choose on sublattice $A$ the representation
-$\hat{S}_{\bm{i}}^{z}=S-\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\,,\quad\hat{S}_{\bm{i}}^{+}=\hat{S}_{\bm{i}}^{x}+i\hat{S}_{\bm{i}}^{y}=\sqrt{2S}\hat{b}_{\bm{i}}^{\phantom{\dagger}}$ (8)
-and on sublattice $B$
+$\hat{S}_{\bm{i}}^{z}=S-\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\,,\quad\hat{S}_{\bm{i}}^{+}=\hat{S}_{\bm{i}}^{x}+i\hat{S}_{\bm{i}}^{y}=\sqrt{2S}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\,,$ (8)
+and on sublattice $B$,
 $\hat{S}_{\bm{i}}^{z}=\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}-S\,,\quad\hat{S}_{\bm{i}}^{+}=\sqrt{2S}\hat{b}_{\bm{i}}^{\dagger}\,.$ (9)
 The correction to the classically expected Néel state is given by the expectation value of the bosonic occupation number $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$.
-In Appendix B we sketch the calculation of the site-dependent correction.
+In Appendix A we sketch the calculation of the site-dependent correction.
 III.3 Quantum Monte Carlo
-In order to acquire deeper information about the electronic correlations on the various lattice geometries and support the mean-field and spin wave data, we employ QMC studies as our secondary approach.
+To acquire deeper information about the electronic correlations on the various lattice geometries and support the mean-field and spin-wave data, we employ QMC studies as our secondary approach.
 In the weak-coupling limit, we use a finite-temperature version of the auxiliary-field QMC (AFQMC) method to simulate the Hubbard model [58, 48, 59].
@@
 $\chi_{O}=\frac{\beta}{N_{s}}\sum_{\bm{i},\bm{j}}^{N_{s}}\left(\big{\langle}\hat{O}_{\bm{i}}\hat{O}_{\bm{j}}\big{\rangle}-\big{\langle}\hat{O}_{\bm{i}}\big{\rangle}\big{\langle}\hat{O}_{\bm{j}}\big{\rangle}\right)$ (10)
-at nonzero Hubbard-$U$ with $\hat{O}_{\bm{i}}$ being either density $\hat{n}_{\bm{i}}=\sum_{\sigma}\hat{n}_{\bm{i},\sigma}$ or spin operators $\hat{S}_{\bm{i}}^{z}$.
+at nonzero Hubbard $U$ with $\hat{O}_{\bm{i}}$ being either density $\hat{n}_{\bm{i}}=\sum_{\sigma}\hat{n}_{\bm{i},\sigma}$ or spin operators $\hat{S}_{\bm{i}}^{z}$.
 To directly compare the QMC results with mean-field, we make use of the bipartite nature of the chosen lattices by constructing an AFM order parameter as
@@
 This approach allows us to access larger lattice sizes compared to the AFQMC method, since SSE simulations scale as $\mathcal{O}(\beta N_{s})$ as opposed to $\mathcal{O}(\beta N_{s}^{3})$ [62].
-The SSE method is based on the Taylor expansion of the partition function $Z$ in some complete basis $\{\left|\alpha\right>\}$ (here we choose the $S^{z}$-basis)
+The SSE method is based on the Taylor expansion of the partition function $Z$ in some complete basis $\{\left|\alpha\right>\}$ (here we choose the $S^{z}$ basis):
 $Z=\sum_{\alpha}\sum_{n=0}^{\infty}\sum_{S_{n}}\frac{(-\beta)^{n}}{n!}\left<\alpha\right|\prod_{i=1}^{n}H_{a_{i},b_{i}}\left|\alpha\right>\,,$ (12)
-$n$ is the current expansion order, $S_{n}$ specifies the operator string $[(a_{1},b_{1}),\ldots,(a_{n},b_{n})]$ and $H_{a_{i},b_{i}}$ is a bond term of the Hamiltonian which operates on bond $b_{i}$.
+where $n$ is the current expansion order, $S_{n}$ specifies the operator string $[(a_{1},b_{1}),\ldots,(a_{n},b_{n})]$, and $H_{a_{i},b_{i}}$ is a bond term of the Hamiltonian which operates on bond $b_{i}$.
 This action can be diagonal or off-diagonal depending on $a_{i}$.
@@
 (11)] rely on the calculation of equal-time two-point spin correlations $\big{\langle}\hat{S}^{z}_{\bm{i}}\hat{S}^{z}_{\bm{j}}\big{\rangle}$.
-As these operators are diagonal in the $S^{z}$-basis, their evaluation within the SSE formalism is trivial.
-Figure 2: (Ia)-(Ic) DOS $\rho(E)$ as a function of the energy $E$ at $\delta=0.01$ for the largest considered lattice sizes ($n_{l}=5$ for the $\{10,3\}$ lattice, $n_{l}=6$ for $\{8,3\}$ and $n_{l}=4$ for $\{8,8\}$, see also Table 1) and (IIa)-(IIc) uniform spin susceptibility $\chi_{S}$ as a function of the temperature $T$ and the number of layers $n_{l}$ in the non-interacting case $U=0$.
+As these operators are diagonal in the $S^{z}$ basis, their evaluation within the SSE formalism is trivial.
+Figure 2: (Ia)–(Ic) Total DOS $\rho(E)$ as a function of the energy $E$ at $\delta=0.01$ for the largest considered lattice sizes ($n_{l}=5$ for the $\{10,3\}$ lattice, $n_{l}=6$ for $\{8,3\}$ and $n_{l}=4$ for $\{8,8\}$, see also Table 1).
+(IIa)–(IIc) Uniform spin susceptibility $\chi_{S}$ and (IIIa)–(IIIc) staggered spin susceptibility $\chi^{x}(Q)$ as a function of the temperature $T$ and the number of layers $n_{l}$ in the noninteracting case $U=0$.
 IV Results
-In the following we study the SU(2) spin symmetric Hubbard and Heisenberg models on three different hyperbolic lattice geometries at the particle-hole symmetric point.
-As shown in Appendix C particle-hole symmetry leads to a logarithmic divergence of the staggered spin susceptibility, provided that the DOS is finite at the Fermi energy.
+In the following we study the SU(2) spin-symmetric Hubbard and Heisenberg models on three different hyperbolic lattice geometries at the particle-hole symmetric point.
+As shown in Appendix B particle-hole symmetry leads to a logarithmic divergence of the staggered spin susceptibility, provided that the DOS is finite at the Fermi energy.
 Hence, the first question that we will address in Sec.
-IV.1 is the DOS in the non-interacting limit.
+IV.1 is the DOS in the noninteracting limit.
 In Sec.
@@
 In Sec.
-IV.3, we then systematically take into account fluctuations with a spin wave approximation for the Heisenberg model, and numerically exact QMC simulations for both models.
+IV.3, we then systematically take into account fluctuations with a spin-wave approximation for the Heisenberg model, and numerically exact QMC simulations for both models.
 IV.1 Non-interacting limit $U=0$
-First we consider the DOS of the $\{10,3\}$ lattice,
+First, we consider the total DOS of the $\{10,3\}$ lattice,
 $\rho(E)=-\frac{1}{\pi N_{s}}\sum_{n}\mathrm{Im}G^{R}(n,E)\,.$ (13)
-The retarded Green function is given by $G^{R}(n,E)=\frac{1}{E-E_{n}+i\delta}$ with an infinitesimal $\delta$ and $E_{n}$ are the discrete energy eigenstates of the tight-binding Hamiltonian $\hat{H}_{t}$.
+The retarded Green function is given by $G^{R}(n,E)=\frac{1}{E-E_{n}+i\delta}$ with an infinitesimal $\delta$ and $E_{n}$ are the discrete energy eigenvalues of the tight-binding Hamiltonian $\hat{H}_{t}$.
 Since our Hamiltonian enjoys particle-hole symmetry, the DOS is symmetric around $E=0$ and the chemical potential for half-band filling is pinned to zero for all temperatures.
@@
 […]
-for the free case.
-In the above, $f$ denotes the Fermi-function.
-Both the DOS and the uniform spin susceptibility are plotted in Fig.
-2(Ia) and 2(IIa).
-The data, in the limit of large values of $n_{l}$, is consistent with a DOS $\rho(E)\propto|E|$ and a corresponding uniform spin susceptibility $\chi_{S}\propto T$.
-The DOS close to the Fermi level as well as the uniform spin susceptibility bear similarities with Dirac systems.
-In Figs.
-2(Ib) and 2(IIb) we consider the $\{8,3\}$ lattice.
+for the free case. In Eq. (14), $f$ denotes the Fermi function. Owing to particle-hole symmetry, the staggered spin susceptibility, see Appendix B, reads:
+[…]
+Thereby, both $\chi^{x}(Q)$ and $\chi_{S}$ depend upon the total DOS. If $\rho(E)\propto|E|$, then $\chi_{S}\propto T$ and $\chi^{x}(Q)$ converges to a constant.
+The DOS, uniform and staggered spin susceptibilities are plotted in Figs.
+2(Ia)–2(IIIa) for the $\{10,3\}$ lattice.
+The same quantities are considered in the (b) panels of Fig.
+2 for the $\{8,3\}$ lattice.
+For these lattices our results suggest a Dirac-like total DOS, but finite-size effects remain too dominant for a definite conclusion.
+Size effects are smeared out by temperature.
+Let us first concentrate on the $\{8,3\}$ lattice.
 At even number of layers, there are no zero-energy eigenstates in the spectrum and $\chi_{S}$ decreases exponentially in the low-temperature limit.
 The activation gap can be attributed to the finite-size gap.
-On the other hand, for an odd number of layers exactly two eigenstates at zero energy are present, thereby resulting in a $T^{-1}$ law of the uniform spin susceptibility in the low-temperature limit.
+However, for an odd number of layers exactly two eigenstates at zero energy are present, thereby resulting in a $T^{-1}$ law of the uniform spin susceptibility in the low-temperature limit.
 But as the number of layers is increased the ratio of zero-energy eigenstates to the total number of eigenstates quickly converges to zero and the magnitude of the $T^{-1}$ law vanishes in the large $n_{l}$ limit.
+For the $\{8,3\}$ lattice, we can compare the data for the susceptibilities, Figs.
+2(IIb) and 2(IIIb), for our two biggest lattices, $n_{l}=5,6$ and observe convergence down to $T/t\simeq 0.01$, a pretty cold temperature given the bandwidth of $6t$.
+Extrapolation of the data from $T/t>0.01$ down to $T=0$ suggest a finite staggered spin susceptibility and a vanishing uniform one.
+Very same conclusions can be drawn for the $\{10,3\}$ lattice.
 We note that the slope of the DOS around $E=0$ and of the spin susceptibility is steeper for the $\{8,3\}$ lattice compared to the former lattice.
 In Euclidean space, the slope is given by the inverse of the Fermi velocity.
-The $\{8,8\}$ lattice on the other hand shows flat band features on the considered finite lattices.
+The $\{8,8\}$ lattice, however, shows flat band features on the considered finite lattices.
 A large number of energy eigenstates are located at zero energy [Fig.
-2(Ic)] and the uniform spin susceptibility diverges in the zero temperature limit [Fig.
-2(IIc)].
+2(Ic)] and the uniform $\chi_{S}$ and staggered $\chi^{x}(Q)$ spin susceptibilities diverge in the zero temperature limit [Figs.
+2(IIc) and 2(IIIc)].
+Figure 3: Single-particle spectral function $A_{1,1}(\omega)$ at $\delta=0.01$ and (a) $n_{l}=5$, (b) $n_{l}=6$, and (c) $n_{l}=4$.
 As mentioned previously, the boundary of hyperbolic lattices has an extensive number of sites such that the bulk, Ref.
@@
 (13) can differ.
-Comparing results, we see that the flat band observed in the $\{8,8\}$ lattice is boundary induced.
+In Fig.
+3, we consider the density of states at a lattice site on the first layer, which we label as $n=1$.
+The bulk density of states, is then given by $A_{1,1}(\omega)=-\frac{1}{\pi}\text{Im}G^{R}(n=1,n^{\prime}=1,\omega)$, where $G^{R}(n=1,n^{\prime}=1,\omega)$ is the real-space retarded Green function between points $n$ and $n^{\prime}$ on the lattice.
+As $n_{l}$ grows this quantity will capture the bulk DOS.
+We plot this quantity in Fig.
+3 for the three lattices we consider and on our largest system size.
+As apparent, we can reproduce the results of Ref.
+[27] for the $\{10,3\}$ and $\{8,3\}$ lattices, and confirm that the bulk DOS takes a finite value at the Fermi energy for all considered lattices.
 IV.2 Mean-field approximation
 IV.2.1 Staggered magnetization and metal-insulator transition
-Figure 3: (a)-(c) AFM order parameter $m^{z}$ as a function of the interaction strength $U$.
-The red line in (a) is a linear fit to the $n_{l}=4$ data of the $\{10,3\}$ lattice close to the critical point $U_{c}\approx 0.76t$ and in (b) to the $n_{l}=6$ data close to $U_{c}\approx 0.54t$.
-Insets: local order parameter $m_{\bm{i}}^{z}$ at $U=5.0t$ for (a) $n_{l}=4$, (b) $n_{l}=5$ and (c) $n_{l}=4$.
+Figure 4: (a)–(c) AFM order parameter $m^{z}$ as a function of the interaction strength $U$.
+The red line in panel (a) is a linear fit to the $n_{l}=4$ data of the $\{10,3\}$ lattice close to the critical point $U_{c}\approx 0.76t$ and in panel (b) to the $n_{l}=6$ data close to $U_{c}\approx 0.55t$.
+Insets: local order parameter $m_{\bm{i}}^{z}$ at $U=5.0t$ for (a) $n_{l}=4$, (b) $n_{l}=5$, and (c) $n_{l}=4$.
 Figure 5: (a) AFM order parameter $m^{z}$ as a function of the temperature $T$ for different interaction strengths as obtained from the mean-field approximation.
@@
 4 show the local magnetic moment $m_{\bm{i}}^{z}$, with a nearly spatially uniform modulus $|m_{\bm{i}}^{z}|$, deep in the AFM phase at $U=5t$.
-The behavior of the three lattices in the weak-coupling limit differ strongly as a consequence of the DOS in the non-interacting case.
-In fact in Appendix C we show, that for a finite DOS at the Fermi level and under the assumption of partial particle-hole symmetry an instability towards AFM can be expected even in the absence of translational symmetry.
-For the Dirac-like $\{10,3\}$ lattice the order parameter $m^{z}$ picks up a nonzero value only above a critical interaction strength $U_{c}\approx 0.76t$.
+The behavior of the three lattices in the weak-coupling limit differ strongly as a consequence of the total DOS in the noninteracting case.
+In fact, in Appendix B we show, that for a finite total DOS at the Fermi level and under the assumption of partial particle-hole symmetry an instability toward AFM can be expected even in the absence of translational symmetry.
+Hence for the Dirac-like lattices we would not expect an AFM instability at small values of the interaction.
+For the open hyperbolic lattices and since the bulk DOS is finite at the Fermi energy, one could question this conclusion.
+In Appendix C, we present a calculation to address the instability of the bulk.
+Here, we place a Hubbard-$U$ term only on the first layer and ask the question if in the infinite-size limit order will occur at infinitesimally small couplings.
+The calculation shows that in this limit there is no instability.
+In fact, using the $C_{p}$ symmetry of the lattices, we can define $p$ unit cells each with $N_{s}/p$ orbitals and a $p$-valued $k$-vector.
+In this notation, which we will also use below, the susceptibility associated with introducing a staggered magnetic field on the first layer reads (see also Appendix C):
+[…]
+[…]
+Here the spectral function $A(k,\omega)$ is a matrix of dimension given by the number of orbitals in the unit cell, and the orbital labeled by 1 belongs to the first layer.
+Equation (16) is generic.
+For a single band system, $A_{1,1}(k,\omega)=\delta(\omega-\epsilon(k))$ such that the generic log-divergence is reproduced.
+However, in our case, when the limit $n_{l}\rightarrow\infty$, $A_{1,1}(k,\omega)$ reflects the bulk DOS and takes a constant value at the Fermi energy.
+Under those circumstances the integral does not diverge.
+Hence for the $\{10,3\}$ and $\{8,3\}$ lattices, we do not expect the onset of long ranged magnetic order at small values of $U$.
+For the Dirac-like $\{10,3\}$ lattice the order parameter $m^{z}$ indeed picks up a nonzero value only above a critical interaction strength $U_{c}\approx 0.76t$.
 Below $U_{c}$ the paramagnetic state is energetically favorable.
-Hartree-Fock mean-field theory for the flat honeycomb lattice predicts a linear increase of the order parameter above the critical interaction $m^{z}\propto(U-U_{c})$, i.e.
-a critical exponent of $\beta=1$ [46, 60, 61].
+Hartree-Fock mean-field theory for the flat honeycomb lattice predicts a linear increase of the order parameter above the critical interaction $m^{z}\propto(U-U_{c})$, i.e., a critical exponent of $\beta=1$ [50, 64, 65].
 The order parameter of the $n_{l}=4$ lattice is well described by a linear fit close to the critical point [Fig.
-3(a)].
-The order parameter on the $\{8,3\}$ exhibits clear odd-even effects in $n_{l}$.
-Let us consider the $n_{l}=1$ lattice, that corresponds to the $p$-site ring with $p=8$.
+4(a)].
+The order parameter on the $\{8,3\}$ lattice exhibits clear odd-even effects in $n_{l}$.
+Let us consider the $n_{l}=1$ lattice, which corresponds to the $p$-site ring with $p=8$.
 Since the rotational symmetry of the ring is shared by the $n_{l}>1$ lattices, a symmetry analysis will prove to be very useful.
-The ground state of the $p=8$ site ring is four-fold degenerate:
+The ground state of the $p=8$ site ring is four fold degenerate:
 |±$|\pm,\sigma\rangle=\hat{c}^{\dagger}_{k=\pm\frac{\pi}{2},\sigma}|0\rangle\,.$ (17)
-In the above $|0\rangle$ is the non-degenerate ground state of the 8-site ring occupied with six electrons, and $\hat{c}^{\dagger}_{k,\sigma}=\frac{1}{\sqrt{p}}\sum_{j=1}^{p}e^{ikj}\hat{c}^{\dagger}_{j,\sigma}$.
+In Eq.
+(17), $|0\rangle$ is the nondegenerate ground state of the eight-site ring occupied with six electrons, and $\hat{c}^{\dagger}_{k,\sigma}=\frac{1}{\sqrt{p}}\sum_{j=1}^{p}e^{ikj}\hat{c}^{\dagger}_{j,\sigma}$.
 Under time-reversal symmetry, $\mathcal{T}$, and $C_{8}$ rotations the states transform as
@@
 $\Delta\sum_{\sigma}\sigma\left[|+,\sigma\rangle\langle-,\sigma|+|-,\sigma\rangle\langle+,\sigma|\right]$ (20)
-are allowed and will lift the ground state degeneracy, yielding
+are allowed and will lift the ground-state degeneracy, yielding
 $|\Psi_{0}\rangle=\frac{1}{2}\left(|+,\uparrow\rangle+|-,\uparrow\rangle\right)\otimes\left(|+,\downarrow\rangle-|-,\downarrow\rangle\right)\,.$ (21)
 This state has a finite staggered magnetization $m^{z}$ for any value of $\Delta$, and is symmetric under combined $C_{8}\mathcal{T}$ symmetries.
-The jump in the magnetization observed at $n_{l}=1$ is a consequence of the degeneracy of the non-interacting ground state and the concomitant divergence of the uniform spin susceptibility.
+The jump in the magnetization observed at $n_{l}=1$ is a consequence of the degeneracy of the noninteracting ground state and the concomitant divergence of the uniform spin susceptibility.
 While the symmetry analysis will apply to arbitrary values of $n_{l}$, the degeneracy only occurs for odd values of $n_{l}$.
 Since the degeneracy is intensive, the jump in the staggered magnetization will be suppressed as a function of $n_{l}$.
-For even values of $n_{l}$ the finite size gap leads to a finite value of $U$ beyond which magnetic ordering will occur.
+For even values of $n_{l}$ the finite-size gap leads to a finite value of $U\approx 0.5t-0.6t$ beyond which magnetic ordering will occur.
+For the largest system size in Fig.
+4(b) with $n_{l}=6$, the data can be described by a linear fit with $U_{c}\approx 0.55t$ similar to the $\{10,3\}$ lattice.
 In the large $n_{l}$ limit we conjecture a Dirac-like system.
-The–in comparison to the $\{10,3\}$ lattice lower–critical interaction strength, that lies within the range of $U_{c}\simeq 0.5t-0.6t$, is consistent with the observation of Sec.
+The—in comparison to the $\{10,3\}$ lattice lower—critical interaction strength, which lies within the range of $U_{c}\simeq 0.5t-0.6t$, is consistent with the observation of Sec.
 IV.1, that the slope of the DOS around $E=0$ and of the uniform spin susceptibility of the $\{8,3\}$ lattice is steeper than the one of the $\{10,3\}$ lattice.
@@
 In Fig.
-4(a) we present the AFM order parameter as a function of temperature $T$, and Fig.
-4(b) the critical temperature as a function of the interaction strength.
+5(a) we present the AFM order parameter as a function of temperature $T$, and in Fig.
+5(b) we present the critical temperature as a function of the interaction strength.
 For the Dirac-like systems our data is consistent with $m^{z}\propto(T-T_{c})^{1/2}$, in the vicinity of $T_{c}$ and $T_{c}\propto(U-U_{c})$, in the vicinity of $U_{c}$.
 These forms are precisely the same as obtained for Dirac systems on Euclidean lattices.
-Hence at the mean-field level the semi-metal to insulator transition in hyperbolic and flat spaces are identical.
-In contrast the flat-band-like $\{8,8\}$ lattice immediately supports AFM ordering for any finite $U$ due to the large DOS at the Fermi level in the non-interacting limit.
-The singularity at $U=0$ for all considered values of $n_{l}$ can be traced back to the divergence of the uniform spin susceptibility and accompanying extensive ground state degeneracy.
+Hence at the mean-field level, the transition in hyperbolic space and the semimetal to insulator transition in flat space are identical.
+In contrast the flat-band-like $\{8,8\}$ lattice immediately supports AFM ordering for any finite $U$ due to the large DOS at the Fermi level in the noninteracting limit.
+The singularity at $U=0$ for all considered values of $n_{l}$ can be traced back to the divergence of the uniform spin susceptibility and accompanying extensive ground-state degeneracy.
 Although the $T=0$ magnetization is singular at $U=0$ the Néel transition temperature seems to follow a powerlaw [Fig.
-4)] as for other flat band systems [62].
-In particular $\lim_{T\rightarrow 0}\lim_{U\rightarrow 0}m^{z}=0$, but $\lim_{U\rightarrow 0}\lim_{T\rightarrow 0}m^{z}>0$.
-Figure 5: (a) Energy eigenvalues $E_{n}$ as obtained from the self-consistent solution of the mean-field Hamiltonian (4), inset: close-up of the Fermi level at $E_{n}=0$, (b) absolute value of the wave function of zero mode $|\psi_{n=1444}(\bm{i})|^{2}$ at $U=0.001t$ and $\sigma=\uparrow$ and (c) magnetization $m^{z}_{\bm{i}}$ at $U=0.4t$ for the $\{8,3\}$ lattice with $n_{l}=5$ layers.
+5(b)] as for other flat band systems [66].
+In particular, $\lim_{T\rightarrow 0}\lim_{U\rightarrow 0}m^{z}=0$, but $\lim_{U\rightarrow 0}\lim_{T\rightarrow 0}m^{z}>0$.
+Figure 6: (a) Energy eigenvalues $E_{n}$ as obtained from the self-consistent solution of the mean-field Hamiltonian (4); inset: close-up of the Fermi level at $E_{n}=0$.
+(b) Absolute value of the wave function of zero mode $|\psi_{n=1444}(\bm{i})|^{2}$ at $U=0.001t$ and $\sigma=\uparrow$ and (c) magnetization $m^{z}_{\bm{i}}$ at $U=0.4t$ for the $\{8,3\}$ lattice with $n_{l}=5$ layers.
 Figure 7: Local magnetization $m^{z}_{\bm{i}}$ as obtained from mean-field approximation as a function of the angle $\phi$ at two different interaction strengths (a) $U=0.4t$ and (b) $U=5t$ for the $\{8,3\}$ lattice with $n_{l}=3$.
-For clarity all sites of one layer are in a separate plot [(I) layer 1, (II) layer 2 and (III) layer 3].
+For clarity all sites of one layer are in a separate plot [(I) layer 1, (II) layer 2, and (III) layer 3].
 IV.2.2 Global antiferromagnetism
-Using the $\{8,3\}$ lattice as an example, we now concentrate on the weak-coupling limit, meaning that the interaction is (much) smaller than the electronic bandwidth in the non-interacting limit.
-Fig.
-5(a) shows the energy eigenvalues of the self-consistent solution of the mean-field Hamiltonian (4) for the $\{8,3\}$ lattice with $n_{l}=5$ layers and different interaction strengths $U$.
-In the non-interacting limit the spectrum is gapless with exactly two zero-energy eigenstates.
+Using the $\{8,3\}$ lattice as an example, we now concentrate on the weak-coupling limit, meaning that the interaction is (much) smaller than the electronic bandwidth in the noninteracting limit.
+Figure 6(a) shows the energy eigenvalues of the self-consistent solution of the mean-field Hamiltonian (4) for the $\{8,3\}$ lattice with $n_{l}=5$ layers and different interaction strengths $U$.
+In the noninteracting limit the spectrum is gapless with exactly two zero-energy eigenstates.
 As for the $n_{l}=1$ case, the wave function of a single zero mode, depicted in Fig.
-5(b), shows broken $C_{8}$ symmetry.
+6(b), shows broken $C_{8}$ symmetry.
 Turning on a finite interaction opens a gap, [inset of Fig.
-5(a)] and the AFM order parameter acquires a finite value.
-As for the $n_{l}=1$ case, a small interaction strength of e.g.
-$U=0.4t$ affects mainly the two eigenstates at the Fermi level [inset of Fig.
-5(a)].
-Figure 5(b) shows the absolute value of the wave function $|\psi_{n}(\bm{i})|^{2}$ of one zero mode in real-space.
-In order to lift the degeneracy of the zero modes and have a uniquely defined pattern of the wave function, we choose a small interaction strength of $U=0.001t$ in Fig.
-5(b).
+6(a)] and the AFM order parameter acquires a finite value.
+As for the $n_{l}=1$ case, a small interaction strength of, e.g., $U=0.4t$ affects mainly the two eigenstates at the Fermi level [inset of Fig.
+6(a)].
+Figure 6(b) shows the absolute value of the wave function $|\psi_{n}(\bm{i})|^{2}$ of one zero mode in real space.
+To lift the degeneracy of the zero modes and have a uniquely defined pattern of the wave function, we choose a small interaction strength of $U=0.001t$ in Fig.
+6(b).
 The wave function has support only on a small number of sites and it breaks the $C_{p=8}$ symmetry of the lattice down to $C_{p/2=4}$.
@@
 The spatial distribution of the order parameter $m^{z}_{\bm{i}}$ follows the pattern of these wave functions [Fig.
-5(c)].
-In particular, at finite $U$ the ground state is unique, breaks time-reversal and $C_{8}$ symmetries, but satisfies:
+6(c)].
+In particular, at finite $U$ the ground state is unique, breaks time-reversal and $C_{8}$ symmetries, but it satisfies:
 $C_{8}\mathcal{T}|\Psi_{0}\rangle=|\Psi_{0}\rangle\,.$ (22)
 For the $n_{l}=1$ case, this just results in AFM, since the $C_{8}$ transformation amounts to a unit translation along the chain.
-For $n_{l}>1$, new magnetic structures emerge, that we will identify with global AFM in these hyperbolic lattices as seen e.g.
-in strained graphene [63].
-First of all, the total magnetic moment vanishes.
+For $n_{l}>1$, new magnetic structures emerge, which we will identify with global AFM in these hyperbolic lattices as seen, e.g., in strained graphene [67].
+First, the total magnetic moment vanishes.
 However, the vanishing of this moment involves cancellation of extensive ferromagnetic moments on each $p$-th sector of the lattice.
 This is depicted in Fig.
-6, where the local magnetization $m_{\bm{i}}^{z}$ is plotted as a function of the angle $\phi$ in the different layers, both in the weak $U=0.4t$ and strong-coupling limit $U=5t$ for the $\{8,3\}$ lattice with $n_{l}=3$ layers.
+7, where the local magnetization $m_{\bm{i}}^{z}$ is plotted as a function of the angle $\phi$ in the different layers, both in the weak $U=0.4t$ and strong-coupling limit $U=5t$ for the $\{8,3\}$ lattice with $n_{l}=3$ layers.
 In the strong-coupling limit the magnetization is almost uniform on the whole lattice.
 At weak coupling the magnetization varies strongly with the angle.
-Summing over all sites included in $\phi\in[0,\pi/4[$ produces a net positive result, i.e.
-a ferromagnetic moment that cancels with the contribution of the sites included in $\phi\in[\pi/4,\pi/2[$.
+Summing over all sites included in $\phi\in[0,\pi/4[$ produces a net positive result, i.e., a ferromagnetic moment that cancels with the contribution of the sites included in $\phi\in[\pi/4,\pi/2[$.
 The effect is very prominent for the $\{8,3\}$ lattice with $n_{l}=5$ layers [Fig.
-5(c)], but we observed similar effects in other lattice geometries as well, also in the absence of zero modes [see Appendix D.1 for more examples].
+6(c)], but we observed similar effects in other lattice geometries as well, also in the absence of zero modes (see Appendix D.1 for more examples).
 It is tempting to identify our magnetic system as an altermagnet [68].
@@
 In this notation, the mean-field Hamiltonian reduces to:
-$\hat{H}_{MF}=$ $\sum_{i,i^{\prime},n,n^{\prime},\sigma}\hat{c}^{\dagger}_{i,n,\sigma}T_{n,n}(i-i^{\prime})\hat{c}^{\phantom{\dagger}}_{i^{\prime},n^{\prime},\sigma}$
+$\hat{H}_{\mathrm{MF}}=$ $\sum_{i,i^{\prime},n,n^{\prime},\sigma}\hat{c}^{\dagger}_{i,n,\sigma}T_{n,n^{\prime}}(i-i^{\prime})\hat{c}^{\phantom{\dagger}}_{i^{\prime},n^{\prime},\sigma}$
 $+$ $\sum_{i,n,n^{\prime}\sigma}\sigma(-1)^{i}\hat{c}^{\dagger}_{i,n,\sigma}M_{n,n^{\prime}}\hat{c}^{\phantom{\dagger}}_{i,n^{\prime},\sigma}\,.$ (23)
-In the above $M$ is a diagonal matrix that encodes the site magnetization in the unit cell. The factor $(-1)^{i}$ accounts for the global AFM. We can now carry out a Fourier transformation to obtain:
-$\hat{H}_{MF}=\sum_{k\in\mathrm{MBZ},\sigma}$ $\left(\hat{\bm{c}}^{\dagger}_{k,\sigma},\hat{\bm{c}}^{\dagger}_{k+Q,\sigma}\right)$ (22)
+In Eq.
+(IV.2.2), $M$ is a diagonal matrix that encodes the site magnetization in the unit cell.
+The factor $(-1)^{i}$ accounts for the global AFM.
+We can now carry out a Fourier transformation to obtain:
+$\hat{H}_{\mathrm{MF}}=\sum_{k\in\mathrm{MBZ},\sigma}$ $\left(\hat{\bm{c}}^{\dagger}_{k,\sigma},\hat{\bm{c}}^{\dagger}_{k+Q,\sigma}\right)$ (24)
 $\begin{pmatrix}T(k)&\sigma M\\ \sigma M^{\dagger}&T(k+Q)\end{pmatrix}\begin{pmatrix}\hat{\bm{c}}^{\phantom{\dagger}}_{k,\sigma}\\ \hat{\bm{c}}^{\phantom{\dagger}}_{k+Q,\sigma}\end{pmatrix}\,.$
-In the above $Q$ is the one-dimensional AFM wave vector, MBZ refers to the magnetic Brillouin zone and $\hat{\bm{c}}^{\dagger}_{k,\sigma}$ is a spinor accounting for the orbital degrees of freedom.
+In Eq.
+(24), $Q$ is the one-dimensional AFM wave vector, MBZ refers to the magnetic Brillouin zone, and $\hat{\bm{c}}^{\dagger}_{k,\sigma}$ is a spinor accounting for the orbital degrees of freedom.
 Using the determinant identity based on the Schur complement, one will readily show that the energy spectrum does not depend on the spin index.
 As such, we cannot understand the observed global AFM in terms of an altermagnet.
-Figure 7: (a) Average correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{b}}$ of the $n_{l}-1$ inner layers and (b) maximal correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{max}}$ as a function of the system size $\frac{1}{\sqrt{N_{s}}}$ at different strengths $\epsilon$ of the infinitesimal staggered field for the (I) $\{10,3\}$, (II) $\{8,3\}$ and (III) $\{8,8\}$ lattice.
-Figure 8: Local correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to classical Néel state as obtained from spin wave approximation for the $\{8,3\}$ lattice with $n_{l}=4$ at $\epsilon=0.01$.
+Figure 8: (a) Average correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{b}}$ of the $n_{l}-1$ inner layers and (b) maximal correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{max}}$ as a function of the system size $\frac{1}{\sqrt{N_{s}}}$ at different strengths $\epsilon$ of the infinitesimal staggered field for the (I) $\{10,3\}$, (II) $\{8,3\}$, and (III) $\{8,8\}$ lattice.
+Figure 9: Local correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to classical Néel state as obtained from spin-wave approximation for the $\{8,3\}$ lattice with $n_{l}=4$ at $\epsilon=0.01$.
 IV.3 Beyond the mean-field approximation
 We now consider approximate as well as numerically exact methods that take into account fluctuations.
-We will first start with the spin wave theory of the Heisenberg model on hyperbolic lattices, and then use Monte Carlo methods for both the Hubbard and Heisenberg models.
-Figure 9: (Ia)-(Ic) Finite-size scaling of the square of the AFM magnetization (order parameter), defined in Eq.
+We will first start with the spin-wave theory of the Heisenberg model on hyperbolic lattices, and then use Monte Carlo methods for both the Hubbard and Heisenberg models.
+Figure 10: (Ia)–(Ic) Finite-size scaling of the square of the AFM magnetization (order parameter), defined in Eq.
 (11), at $\beta t=100$ for different values of $U/t$ on the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice respectively (from left to right).
@@
 The data points for $U\to\infty$ refer to the Heisenberg model at $\beta J=100$, calculated by SSE.
-(IIa)-(IIc) Comparison of the squared magnetization, the squared magnetization of the bulk $\tilde{m}_{\textrm{AFM}}^{2}$ and the spin-spin correlation $\langle S_{\bm{r}}^{z}S_{0}^{z}\rangle$ of two sites with the largest distance to each other on the respective lattice, where $S_{0}^{z}$ is the site closest to $\phi=0$ (for the $\{10,3\}$-lattice this is negative since opposing sites lie on different sublattices).
+(IIa)–(IIc) Comparison of the squared magnetization, the squared magnetization of the bulk $\tilde{m}_{\textrm{AFM}}^{2}$ and the spin-spin correlation $\langle S_{\bm{r}}^{z}S_{0}^{z}\rangle$ of two sites with the largest distance to each other on the respective lattice, where $S_{0}^{z}$ is the site closest to $\phi=0$ (for the $\{10,3\}$ lattice this is negative since opposing sites lie on different sublattices).
 Here, we use $U/t=6$.
@@
 [69, 70], where spin-wave calculations were formulated for spin glasses.
-The method is summarized in Appendix B.
+The method is summarized in Appendix A.
 To avoid zero modes, we include a small staggered field of magnitude $\epsilon$ and take the limit $\epsilon\rightarrow 0$.
@@
 The left panels show the average correction within the bulk $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{b}}$, where we defined the bulk as the $n_{l}-1$ inner layers, and the right panels show the maximal correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{max}}$, which can always be found on the sites closest to the angle $\phi=0$ in the outermost layer and all sites, that can be obtained by performing symmetry-allowed rotations.
-As described in Appendix B the limit of $\epsilon\rightarrow 0$ and the thermodynamic limit have to be taken carefully.
-Taking the limit $\epsilon\rightarrow 0$ on finite lattices leads to divergencies in the correction [Fig.
-7] and only in the thermodynamic limit $\frac{1}{\sqrt{N_{s}}}\rightarrow 0$ the correct results are recovered.
+As described in Appendix A the limit of $\epsilon\rightarrow 0$ and the thermodynamic limit have to be taken carefully.
+Taking the limit $\epsilon\rightarrow 0$ on finite lattices leads to divergencies in the correction (Fig.
+8) and only in the thermodynamic limit $\frac{1}{\sqrt{N_{s}}}\rightarrow 0$ the correct results are recovered.
 The average corrections in the bulk converge to values smaller than $\frac{1}{2}$ on the largest available lattice sizes for all three lattice geometries [Fig.
-7(a)], meaning that the Néel state in the bulk is weakened, but still present in the case of spin $1/2$ degrees of freedom.
-The maximal correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{max}}$ on the other hand side is above $\frac{1}{2}$, such that it is large enough to destroy the magnetic moment on the corresponding sites.
+8(a)], meaning that the Néel state in the bulk is weakened, but still present in the case of spin $1/2$ degrees of freedom.
+The maximal correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle_{\mathrm{max}}$, however, is above $\frac{1}{2}$, such that it is large enough to destroy the magnetic moment on the corresponding sites.
 In Fig.
@@
 Therefore the correction on a given site is larger the greater the distance to the nearest site with $q=3$ neighbors.
-This result is consistent with the absence of long-range order in one dimension, as predicted by spin wave theory.
+This result is consistent with the absence of long-range order in one dimension, as predicted by spin-wave theory.
 Similar qualitative observations for the other lattice geometries can be found in Fig.
 17 of Appendix D.2.
-Figure 10: SSE calculated squared magnetization of the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice respectively (from left to right), at $\beta J=100$, together with the bulk and edge contribution.
+Figure 11: SSE calculated squared magnetization of the $\{10,3\}$, $\{8,3\}$, and $\{8,8\}$ lattice, respectively (from left to right), at $\beta J=100$, together with the bulk and edge contribution.
 The data points were fitted with a second order polynomial.
@@
 As reference point $S_{0}^{z}$ we chose the site closest to $\phi=0$.
-Figure 12: QMC data for the (Ia)-(Ic) uniform spin susceptibility $\chi_{S}$ and (IIa)-(IIc) charge susceptibility $\chi_{C}$ at $U=5t$ on the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice respectively (from left to right).
+Figure 13: QMC data for the (Ia)–(Ic) uniform spin susceptibility $\chi_{S}$ and (IIa)–(IIc) charge susceptibility $\chi_{C}$ at $U=5t$ on the $\{10,3\}$, $\{8,3\}$, and $\{8,8\}$ lattice, respectively (from left to right).
 IV.3.2 Quantum Monte Carlo
 The aim of this section is to use numerically exact approaches, AFQMC and SSE, to check the validity of the above statements.
-Figures 9(Ia)-9(Ic) plot the staggered spin-spin-correlations as defined in Eq.
+Figures 10(Ia)–10(Ic) plot the staggered spin-spin correlations as defined in Eq.
 (11) as a function of $\frac{1}{\sqrt{N_{s}}}$.
@@
 First, since the number of sites grow rapidly with $n_{l}$ we only have a limited set of data.
-Second, we measure the square of the order parameter such that when the staggered magnetization is small (i.e.
-at weak coupling or in the proximity of a quantum phase transition) very large lattices are required to obtain reliable results.
-We note that a pinning field approach [48] that was used for the Hubbard model on the honeycomb lattice may be a alternative and promising method for future investigations.
+Second, we measure the square of the order parameter such that when the staggered magnetization is small (i.e., at weak coupling or in the proximity of a quantum phase transition) very large lattices are required to obtain reliable results.
+We note that a pinning field approach [52] that was used for the Hubbard model on the honeycomb lattice may be an alternative and promising method for future investigations.
 For the $\{10,3\}$ and $\{8,3\}$ lattices [Figs.
-9(Ia) and 9(Ib)] our data at $\beta t=100$ ($\beta J=100$) is representative of the ground-state properties for the considered values of $n_{l}$.
+10(Ia) and 10(Ib)] our data at $\beta t=100$ ($\beta J=100$) is representative of the ground-state properties for the considered values of $n_{l}$.
+In the strong-coupling limit, the results of both lattices are consistent with AFM long range ordering.
 For the $\{8,3\}$ lattice ordering sets in at $U_{c}\simeq 2t$.
@@
 Hence, the trends in the value of $U_{c}$, $(U_{c}^{\{10,3\}}>U_{c}^{\{8,3\}}$) between the two lattices match our mean-field analysis.
-It is beyond the scope of this article to determine critical exponent of the order parameter: $m\simeq(U-U_{c})^{\beta}$.
+It is beyond the scope of this article to determine the critical exponent of the order parameter: $m\simeq(U-U_{c})^{\beta}$.
 In Figs.
-9(Ia) and 9(Ib), we equally consider the Heisenberg model.
+10(Ia) and 10(Ib), we equally consider the Heisenberg model.
 For this model, the SSE allows us to reach much larger lattices such that the extrapolation is more robust.
 In Fig.
-9(Ic) we consider the flat band system $\{8,8\}$.
+10(Ic) we consider the flat band system $\{8,8\}$.
 With the fermion QMC we can only access three system sizes for the extrapolation.
@@
 In Figs.
-9(IIa)-9(IIc), we also present the AFM magnetization of the bulk $\tilde{m}^{2}_{AFM}$ by only summing over sites of the $n_{l}-1$ inner layers.
-The bulk magnetization is larger than the total magnetization including the edge, consistent with our results from spin wave theory.
+10(IIa)–10(IIc), we also present the AFM magnetization of the bulk $\tilde{m}^{2}_{\mathrm{AFM}}$ by only summing over sites of the $n_{l}-1$ inner layers.
+The bulk magnetization is larger than the total magnetization including the edge, consistent with our results from spin-wave theory.
 This becomes more obvious when we consider the bulk and edge contribution to the magnetization calculated by SSE, as presented in Fig.
-10.
+11.
 This method allows a more accurate study of the magnetization in the strong-coupling limit as larger lattices can be simulated.
 While the edge magnetization is slightly lower than the total magnetization, the bulk contribution is larger by a factor of approximately $\sqrt{2}$ for all considered lattices.
-This specifies our findings from spin wave theory that fluctuations are enhanced on the boundary.
+This specifies our findings from spin-wave theory that fluctuations are enhanced on the boundary.
 The SSE calculations suggest more precisely, that there is still long range order on the edge.
-Finally, we note that due to the sub-extensiveness of the bulk, the magnetization is determined by the edge contribution.
+Finally, we note that due to the subextensiveness of the bulk, the magnetization is determined by the edge contribution.
 This is supported by the SSE data of Fig.
@@
 To compare the results from mean-field with QMC, we present spin correlations plotted in the Poincaré disk of the $\{10,3\}$ lattice as shown in Fig.
-11.
+12.
 Again considering the weak-coupling regime, similar patterns of alternating positive and negative correlations can be observed in the third layer.
 This can be directly compared to the mean-field results of the order parameter for the $\{10,3\}$ lattice in Fig.
-13(a), see Appendix.
+14(a); see Appendix.
 For the presentation of the QMC results, we choose the $\{10,3\}$ lattice, since it offers the clearest visualization of the observed effect.
@@
 Nevertheless, the observed spin fluctuations support the notion that in the symmetry-broken phase at large $U/t$ the $C_{10}$ and time-reversal symmetry are broken, but $C_{10}\mathcal{T}$ remains a symmetry of the state.
-In order to support the mean-field observation that the system is in an AFM Mott insulating state for strong electronic correlations, we present the spin and charge susceptibility for $U=5t$ on the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice in Figs.
-12(Ia)-12(Ic) and 12(IIa)-12(IIc) respectively.
+To support the mean-field observation that the system is in an AFM Mott insulating state for strong electronic correlations, we present the spin and charge susceptibilities for $U=5t$ on the $\{10,3\}$, $\{8,3\}$, and $\{8,8\}$ lattices in Figs.
+13(Ia)–13(Ic) and 13(IIa)–13(IIc) respectively.
 As expected for an AFM state spin waves lead to a spin susceptibility $\chi_{S}$ that converges to a finite value as $T\to 0$.
 The data shows that there is a buildup of low energy states as one increases $n_{l}$.
-On the other hand, the charge susceptibility $\chi_{C}$ decays exponentially, as expected for a charged, gaped system.
-Figures 12(IIa)-12(IIc) show that as $n_{l}$ grows the charge susceptibility remains activated in the low-temperature limit.
+However, the charge susceptibility $\chi_{C}$ decays exponentially, as expected for a charged, gaped system.
+Figures 13(IIa)–13(IIc) show that as $n_{l}$ grows the charge susceptibility remains activated in the low-temperature limit.
 V Discussion and Conclusion
@@
 Owing to a theorem by Lieb [55], the ground state on a finite lattice is a spin singlet.
-The lowest energy state in each total spin sector builds the Anderson tower of states [67], that collapses in the thermodynamic limit.
+The lowest energy state in each total spin sector builds the Anderson tower of states [71], which collapses in the thermodynamic limit.
 This is the mechanism that allows for the broken-symmetry quantum AFM ground state without violating the aforementioned Lieb theorem in dimensions greater or equal to two.
@@
 As a consequence, the very notion of dimensionality and role of fluctuations can be questioned.
-While our mean-field calculations for the Hubbard model show long-range order on the boundary and in the bulk, a spin wave calculation for the Heisenberg model shows that fluctuations inhibit order at the boundary.
-As demonstrated by our QMC simulations for both the strong-coupling Hubbard and Heisenberg models, this turns out to be an artifact of the spin wave approximation.
-It is known that spin waves fail to predict that weakly coupling chains with non-frustrating interactions is a relevant perturbation [68] that leads to long range magnetic order.
+While our mean-field calculations for the Hubbard model show long-range order on the boundary and in the bulk, a spin-wave calculation for the Heisenberg model shows that fluctuations inhibit order at the boundary.
+As demonstrated by our QMC simulations for both the strong-coupling Hubbard and Heisenberg models, this turns out to be an artifact of the spin-wave approximation.
+It is known that spin waves fail to predict that weakly coupling chains with nonfrustrating interactions is a relevant perturbation [72] that leads to long range magnetic order.
 However, even if the ground state is ordered, it would be of interest to investigate the spin dynamical response on the boundary, with the aim to assess whether proximity to one-dimensional physics is visible at high frequencies as in Ref.
@@
 It is known that weak frustrating couplings between chains do not necessarily lead to magnetic ordering [73], thereby offering the possibility of realizing different phases in the bulk and on the boundary of hyperbolic lattices.
-As for the honeycomb lattice, the Dirac-like character of the DOS close to the Fermi level of the $\{10,3\}$ and $\{8,3\}$ lattices lead to a finite critical interaction strength $U_{c}$ for the onset of AFM ordering.
+For the considered finite lattices with open boundary conditions, the $\{8,8\}$ lattice shows similarities to flat-band systems, with a high total DOS at the Fermi level, which results in the presence of AFM ordering at any finite interaction $U$.
+The $\{8,3\}$ lattice displays odd-even effects with a vanishing (finite) DOS at the Fermi level for an even (odd) number of layers.
+As for the honeycomb lattice, the Dirac-like character of the DOS close to the Fermi level of the $\{10,3\}$ and $\{8,3\}$ lattices in the limit of large $n_{l}$ leads to a finite critical interaction strength $U_{c}$ for the onset of AFM ordering.
 At the mean-field level the order parameter exponent takes the value $\beta=1$ as for the Euclidean case.
-On Euclidean lattices, great progress has been achieved in the understanding of the Gross-Neveu transition, by using renormalization group invariant quantities and finite size scaling to analyze the data [47, 48, 49].
+On Euclidean lattices, great progress has been achieved in the understanding of the Gross-Neveu transition, by using renormalization group invariant quantities and finite-size scaling to analyze the data [51, 52, 53].
 For the hyperbolic lattices our analysis is very rudimentary due to the very difficulty of defining a continuum limit and the accompanying critical theory.
@@
 Nevertheless, our data is consistent with a finite-$U$ Mott transition.
+The observed metal-insulator transition deserves more care since the bulk and total DOS of the noninteracting system differ on hyperbolic lattices with open boundaries.
+To define these quantities precisely, let us consider a subset $M$ of bulk sites of the total number of sites, $N_{s}$.
+The bulk DOS is given by
+$N_{\text{Bulk}}(\omega)=-\frac{1}{\pi V_{M}}\sum_{i\in M}\lim_{N_{s}\rightarrow\infty}\text{Im}G^{R}_{i,i}(\omega)\,,$ (25)
+where $G^{R}_{i,j}(\omega)$ is the real-space retarded Green function between site $i$ and $j$ and $V_{M}$ the number of sites in the set $M$.
+The total DOS is defined as in Eq.
+(13).
+We have shown explicitly that for the Dirac-like lattice, both bulk—with finite DOS at the Fermi energy—and the total system with Dirac-like DOS do not show magnetic instabilities.
+However, systems with periodic boundary conditions, that have a finite DOS at the Fermi level and that are particle-hole symmetric, should show magnetic instabilities and hence no metal-insulator transition.
+In this sense, the observed metal-to-insulator transition is boundary induced.
 The thermodynamic properties of the AFM Mott insulating state are the very same as in the Euclidean case.
@@
 Additionally the magnetic patterns are invariant under a combination of the original rotation group of the lattice and time-reversal symmetry $C_{p}\mathcal{T}$.
-Global AFM occurs e.g.
-also in finite graphene samples with zig-zag edges [70, 71, 72, 73, 74, 75, 76] or in more complex pattern, when the sample is subject to a strain [77, 63, 78].
-The absence of negative sign problem in QMC simulations hinges on symmetries (e.g.
-combined time-reversal and a U(1) conservation) [79, 80] that are not broken by the hyperbolic geometry.
-Hence all models that are amenable to negative sign-free QMC simulations such as SU(N) Hubbard-Heisenberg models [81], SU(N) Kondo lattices [82, 61], and Su-Schrieffer-Heeger Hamiltonians [83, 84] can be studied.
+Global AFM occurs, e.g., also in finite graphene samples with zig-zag edges [74, 75, 76, 77, 78, 79, 80] or in more complex pattern, when the sample is subject to a strain [81, 67, 82].
+The absence of negative sign problem in QMC simulations hinges on symmetries (e.g., combined time-reversal and a U(1) conservation) [83, 84] that are not broken by the hyperbolic geometry.
+Hence all models that are amenable to negative sign-free QMC simulations such as SU(N) Hubbard-Heisenberg models [85], SU(N) Kondo lattices [86, 87], and Su-Schrieffer-Heeger Hamiltonians [88, 89, 90, 91] can be studied.
 It remains to be seen if novel phenomena can be observed due the hyperbolic geometry.
@@
 (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at the Leibniz Supercomputing Centre (www.lrz.de).
-FA and JI thank the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490), GR and AG the DFG funded SFB 1170 on Topological and Correlated Electronics at Surfaces and Interfaces.
-FA acknowledges financial support from the German Research Foundation (DFG) under the grant AS 120/16-1 (Project number 493886309) that is part of the collaborative research project SFB Q-M&S funded by the Austrian Science Fund (FWF) F 86.
-Appendix A Lattice sizes
-Table 1 summarizes the lattices used in the main text and their respective number of lattice sites $N_{s}$ as a function of the number of layers $n_{l}$.
-$\{10,3\}$$\{8,3\}$$\{8,8\}$
-$n_{l}=1$ 10 8 8
-$2$ 80 48 56
-$3$ 490 200 392
-$4$ 2880 768 2744
-$5$ 16810 2888
-$6$ 10800
-Table 1: Number of lattice sites for considered lattice geometries as a function of the number of layers $n_{l}$.
-Appendix B Spin wave approximation
+F.A.
+and J.I.
+thank the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490), G.R.
+and A.G.
+acknowledge the DFG funded SFB 1170 on Topological and Correlated Electronics at Surfaces and Interfaces.
+F.A.
+acknowledges financial support from the German Research Foundation (DFG) under the grant AS 120/16-1 (Project number 493886309) that is part of the collaborative research project SFB Q-M&S funded by the Austrian Science Fund (FWF) F 86.
+Appendix A Spin-wave approximation
 […]
@@
 Here we defined the site-dependent coordination number $z_{\bm{i}}=\sum_{\bm{j}}T_{\bm{i},\bm{j}}$. We can write the equations of motions for the boson operators in a matrix representation as
-$\frac{d}{dt}\hat{\bm{b}}=i\begin{pmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{pmatrix}\hat{\bm{b}}$ (24)
+$\frac{d}{dt}\hat{\bm{b}}=i\begin{pmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{pmatrix}\hat{\bm{b}}\,,$ (27)
 with
-$\hat{\bm{b}}^{\dagger}$ $=$ $(\hat{b}_{\bm{i}=1}^{\dagger},\dots,\hat{b}_{\bm{i}=N}^{\dagger},\hat{b}_{\bm{i}=1}^{\phantom{\dagger}},\dots,\hat{b}_{\bm{i}=N}^{\phantom{\dagger}})\,,$
+$\hat{\bm{b}}^{\dagger}$ $=$ $(\hat{b}_{\bm{i}=1}^{\dagger},\dots,\hat{b}_{\bm{i}=N_{s}}^{\dagger},\hat{b}_{\bm{i}=1}^{\phantom{\dagger}},\dots,\hat{b}_{\bm{i}=N_{s}}^{\phantom{\dagger}})\,,$
 $P_{\bm{i},\bm{j}}$ $=$ $Sz_{\bm{i}}\delta_{\bm{i},\bm{j}}\,,\quad Q_{\bm{i},\bm{j}}=ST_{\bm{i},\bm{j}}\,.$ (28)
-In order to diagonalize the Hamiltonian of Eq. (23) we perform a Bogoliubov transformation
+To diagonalize the Hamiltonian of Eq. (26) we perform a Bogoliubov transformation,
 $\begin{pmatrix}-\Omega&0\\ 0&\Omega\\ \end{pmatrix}=\begin{pmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{pmatrix}^{-1}\begin{pmatrix}-P&-Q\\ \bar{Q}&\bar{P}\\ \end{pmatrix}\begin{pmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{pmatrix}\,.$ (29)
-To ensure, that the new boson operators $\hat{d}_{\bm{i}}^{\dagger}$, that we obtain from $\hat{b}_{\bm{i}}^{\dagger}=\bar{f}_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\phantom{\dagger}}+g_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\dagger}$, still fulfill the canonical commutation relations, we need to impose a condition on the entries of $f$ and $g$
+To ensure that the new boson operators $\hat{d}_{\bm{i}}^{\dagger}$, which we obtain from $\hat{b}_{\bm{i}}^{\dagger}=\bar{f}_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\phantom{\dagger}}+g_{\bm{i},\bm{j}}\hat{d}_{\bm{j}}^{\dagger}$, still fulfill the canonical commutation relations, we need to impose a condition on the entries of $f$ and $g$:
 $\begin{pmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{pmatrix}\begin{pmatrix}-1&0\\ 0&1\\ \end{pmatrix}\begin{pmatrix}\bar{g}&f\\ \bar{f}&g\\ \end{pmatrix}^{\dagger}=\begin{pmatrix}-1&0\\ 0&1\\ \end{pmatrix}\,.$ (30)
@@
 (30) can be satisfied by ensuring, that all eigenvectors of $M$ are orthonormal with respect to the scalar product $\langle\cdot,\cdot\rangle_{B}$.
-If the norm of an eigenvector $\langle\bm{x}_{n},\bm{x}_{n}\rangle_{B}>0$ is nonzero, it can also be shown, that the corresponding eigenvalue is real.
-In order to avoid eigenvectors with a vanishing $B$ norm we add an infinitesimal staggered magnetic field $\hat{H}_{\epsilon}=-\epsilon\sum_{\bm{i}}(-1)^{\bm{i}}\hat{S}_{\bm{i}}^{z}$.
+If the norm of an eigenvector $\langle\bm{x}_{n},\bm{x}_{n}\rangle_{B}>0$ is nonzero, then it can also be shown that the corresponding eigenvalue is real.
+To avoid eigenvectors with a vanishing $B$ norm we add an infinitesimal staggered magnetic field $\hat{H}_{\epsilon}=-\epsilon\sum_{\bm{i}}(-1)^{\bm{i}}\hat{S}_{\bm{i}}^{z}$.
 After the calculation, limits have to be taken carefully, by first taking the thermodynamic limit and then the limit of $\epsilon\rightarrow 0$.
-The site-dependent correction to the classical AFM state can be calculated by the mean boson occupation number at $T=0$
+The site-dependent correction to the classical AFM state can be calculated by the mean boson occupation number at $T=0$:
 $\langle\hat{S}_{\bm{i}}^{z}\rangle=S-\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle=S-\sum_{\bm{j}}|f_{\bm{i},\bm{j}}|^{2}\,.$ (31)
-Appendix C Spin susceptibility
+Appendix B Staggered spin susceptibility
 In this Appendix we show that the assumption of partial particle-hole symmetry and a finite DOS at the Fermi surface, suffices to argue a logarithmic divergence of the staggered spin susceptibility.
@@
 $\hat{P}^{-1}_{\sigma}\hat{c}_{\bm{i},\sigma^{\prime}}^{\dagger}\hat{P}_{\sigma}=\delta_{\sigma,\sigma^{\prime}}(-1)^{\bm{i}}\hat{\gamma}_{\bm{i},\sigma^{\prime}}^{\phantom{\dagger}}+\left(1-\delta_{\sigma,\sigma^{\prime}}\right)\hat{\gamma}_{\bm{i},\sigma^{\prime}}^{\dagger}\,.$ (32)
-Without loss of generality we can consider the staggered spin susceptibility in $x$-direction
+Without loss of generality we can consider the staggered spin susceptibility in the $x$ direction:
 […]
@@
 As apparent, the partial particle-hole transformation maps the staggered susceptibility in the particle-hole channel to the uniform one in the particle-particle channel.
-The latter is nothing but the Cooper instability, that we now show to be present on any graph with finite DOS at the Fermi energy.
+The latter is nothing but the Cooper instability, which we now show to be present on any graph with finite DOS at the Fermi energy.
 To do so, we introduce a canonical transformation $U$, that diagonalizes the Hamiltonian
-$\hat{H}_{t}=\sum_{\bm{i},\bm{j},\sigma}h_{\bm{i},\bm{j}}\hat{\gamma}_{\bm{i},\sigma}^{\dagger}\hat{\gamma}_{\bm{j},\sigma}^{\phantom{\dagger}}=\sum_{x}\lambda_{x}\hat{\eta}_{x,\sigma}^{\dagger}\hat{\eta}_{x,\sigma}^{\phantom{\dagger}}$ (31)
+$\hat{H}_{t}=\sum_{\bm{i},\bm{j},\sigma}h_{\bm{i},\bm{j}}\hat{\gamma}_{\bm{i},\sigma}^{\dagger}\hat{\gamma}_{\bm{j},\sigma}^{\phantom{\dagger}}=\sum_{x}\lambda_{x}\hat{\eta}_{x,\sigma}^{\dagger}\hat{\eta}_{x,\sigma}^{\phantom{\dagger}}\,,$ (34)
 with new fermionic operators $\hat{\eta}_{x,\sigma}^{\dagger}=\sum_{\bm{i}}\hat{\gamma}_{\bm{i},\sigma}^{\dagger}U_{\bm{i},x}$. Applying this transformation yields
@@
 ×$\times\sum_{x,y}f_{x,y}e^{\tau(\lambda_{x}+\lambda_{y})}\langle\hat{\eta}_{x,\uparrow}^{\dagger}\hat{\eta}_{x,\uparrow}^{\phantom{\dagger}}\rangle\langle\hat{\eta}_{y,\downarrow}^{\dagger}\hat{\eta}_{y,\downarrow}^{\phantom{\dagger}}\rangle$
-$=$ $\frac{1}{N_{s}}\sum_{x,y}\frac{f_{x,y}}{\lambda_{x}+\lambda_{y}}\left[\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}+\tanh{\left(\frac{\beta}{2}\lambda_{y}\right)}\right]$
+$=$ $\frac{1}{N_{s}}\sum_{x,y}\frac{f_{x,y}}{\lambda_{x}+\lambda_{y}}\left[\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}+\tanh{\left(\frac{\beta}{2}\lambda_{y}\right)}\right]\,,$
 with $f_{x,y}=\sum_{\bm{i},\bm{j}}(U^{\dagger})_{x,\bm{i}}U_{\bm{j},x}(U^{\dagger})_{y,\bm{i}}U_{\bm{j},y}$.
-If the DOS is finite at zero energy, the Fermi level here, we can show, that the term with $x=y$ has a logarithmic divergency
-$\chi^{x}(Q)$ $=$ $\frac{1}{N_{s}}\sum_{x}\frac{f_{x,x}}{\lambda_{x}}\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}+R$ (33)
+For our specific case, the hopping Hamiltonian is symmetric such that $U$ is an orthogonal matrix.
+As a consequence, $f_{x,y}=\delta_{x,y}$.
+Hence,
+$\chi^{x}(Q)$ $=$ $\frac{1}{N_{s}}\sum_{x}\frac{1}{\lambda_{x}}\tanh{\left(\frac{\beta}{2}\lambda_{x}\right)}$
+[…]
+where the DOS is given by Eq.
+(13).
+The last approximation is in the low-temperature limit and $W$ is a high-energy cutoff.
+Provided that the DOS is finite at the Fermi energy, an instability toward AFM ordering is generic even in the absence of translational symmetry.
+Appendix C Instability of the bulk
+To study the stability of the bulk toward AFM ordering, we place an AFM perturbation $\hat{H}_{I}$ on the inner ring of the lattice
+$\hat{H}$ $=$ $\hat{H}_{t}+\hat{H}_{I}\,,$ (37)
+[…]
+$\hat{H}_{I}$ $=$ $\sum_{i,\sigma}m\sigma(-1)^{i}\hat{c}^{\dagger}_{i,1,\sigma}\hat{c}^{\phantom{\dagger}}_{i,1,\sigma}=\sum_{k,\sigma}m\sigma\hat{c}^{\dagger}_{k,1,\sigma}\hat{c}^{\phantom{\dagger}}_{k+Q,1,\sigma}\,.$
+In Eq.
+(37), we used the notation as introduced in Eq.
+(IV.2.2), where we defined a unit cell $i$ with the help of the $C_{p}$ symmetry of the lattice.
+The orbital $n=1$ refers to the site on the inner layer included in the unit cell $i$.
+Equation (37) is equivalent to Eq.
+(IV.2.2) by setting $M_{1,1}=m$ and $M_{n,n^{\prime}}=0$ otherwise.
+The susceptibility with respect to the perturbation $\hat{H}_{I}$ can be written as
 […]
 […]
-where all other terms in the sum with $x\neq y$ are summarized in the term $R$.
-The last approximation is in the low-temperature limit and $W$ is a high-energy cutoff.
-In principle terms could appear in $R$, that lead to the cancellation of the logarithmic divergency under certain conditions.
-But in general the occurrence of an instability towards AFM ordering is generic even in the absence of translational symmetry.
+[…]
+×$\times A_{1,1}(k,\omega)A_{1,1}(k+Q,\omega^{\prime})\,.$
+Here, we defined the Green function $G_{1,1}(k,\tau)=-\langle\mathcal{T}\hat{c}_{k,1}^{\phantom{\dagger}}(\tau)\hat{c}_{k,1}^{\dagger}\rangle_{0}$ with the imaginary-time ordering operator $\mathcal{T}$ and the expectation value $\langle\dots\rangle_{0}$ with respect to $\hat{H}_{t}$.
+The spectral function $A_{1,1}(k,\omega)=-\frac{1}{\pi}\mathrm{Im}G_{1,1}(k,i\omega_{m}\rightarrow\omega+i\delta)$ is defined via analytical continuation of the Green function with an infinitesimal $\delta$ and the fermionic Matsubara frequencies $\omega_{m}$.
+Using particle-hole symmetry, $A_{1,1}(k,\omega)=A_{1,1}(k+Q,-\omega)$, we can further simplify the last equation to obtain:
+[…]
+×$\times\sum_{k}A_{1,1}(k,\omega)A_{1,1}(k,\omega^{\prime})\,.$
+Since $A_{1,1}(k,\omega)$ takes a finite, constant value at the Fermi energy [27] we can replace the $k$-sum over the spectral functions by a finite number.
+Carrying out the integration leads to a finite value of $\chi(Q,i\Omega_{m}=0)$ such that no bulk magnetic instability is present.
 Appendix D Additional data
-In the following section we present some additional data.
+In the following Appendix we present some additional data.
 D.1 Mean-field approximation
-Figure 13: (I) Local DOS $\rho(E=0,\bm{i})$ at zero energy in the non-interacting limit $U=0$ at $\delta=0.01$ and (II) magnetization $m_{\bm{i}}^{z}$ as obtained from mean-field approximation for the $\{10,3\}$ lattice and interaction strength $U=0.8t$.
+Figure 14: (I) Local DOS $\rho(E=0,\bm{i})$ at zero energy in the noninteracting limit $U=0$ at $\delta=0.01$ and (II) magnetization $m_{\bm{i}}^{z}$ as obtained from mean-field approximation for the $\{10,3\}$ lattice and interaction strength $U=0.8t$.
 Figure 15: Same as in Fig. 14, but for the $\{8,3\}$ lattice and interaction strength (IIa) $U=0.4t$ and (IIb) $U=0.58t$.
@@
 In the upper panels of Figs.
-13-15 we plot the local DOS at zero energy
-$\rho(E,\bm{i})=-\frac{1}{\pi N_{s}}\sum_{n}|U_{\bm{i},n}|^{2}\mathrm{Im}G^{R}(n,E)$ (34)
+1416 we plot the local DOS at zero energy,
+$\rho(E,\bm{i})=-\frac{1}{\pi N_{s}}\sum_{n}|U_{\bm{i},n}|^{2}\mathrm{Im}G^{R}(n,E)\,,$ (40)
 for the different lattice geometries both with $n_{l}=3$ and $n_{l}=4$ layers.
@@
 Here we defined a canonical transformation $U$, that diagonalizes the mean-field Hamiltonian (4).
-They all show a unique pattern of low and high DOS, that is compatible with the $C_{p}$ symmetry of the lattices.
+They all show a unique pattern of low and high DOS, which is compatible with the $C_{p}$ symmetry of the lattices.
 The lower panels of Figs.
-13-15 show the corresponding local order parameter in real-space, where again local ferromagnetic and AFM moments emerge in the different sectors of the lattice.
+1416 show the corresponding local order parameter in real space, where again local ferromagnetic and AFM moments emerge in the different sectors of the lattice.
 But in sum of all lattice sites only the AFM moment remains.
-D.2 Spin wave approximation
-The results for the site-dependent correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to the classical Néel state obtained from the spin wave approximation for the $\{10,3\}$ and $\{8,3\}$ lattice in Figs.
-16(I) and 16(II) are similar to the results for the $\{8,3\}$ lattice described in the main text [Fig.
-8].
+D.2 Spin-wave approximation
+The results for the site-dependent correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to the classical Néel state obtained from the spin-wave approximation for the $\{10,3\}$ and $\{8,3\}$ lattice in Figs.
+17(I) and 17(II) are similar to the results for the $\{8,3\}$ lattice described in the main text [Fig.
+9].
 For the $\{8,8\}$ lattice the results differ slightly due to the lattice geometry.
@@
 The sites on the edge also form short one-dimensional chains, but larger effective interactions are present in comparison to the other two considered lattice geometries.
-Figure 16: Local correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to classical Néel state as obtained from spin wave approximation for the (I) $\{10,3\}$, (II) $\{8,3\}$ and (III) $\{8,8\}$ lattice and varying system size $n_{l}$ at $\epsilon=0.01$.
+Figure 17: Local correction $\langle\hat{b}_{\bm{i}}^{\dagger}\hat{b}_{\bm{i}}^{\phantom{\dagger}}\rangle$ to classical Néel state as obtained from spin-wave approximation for the (I) $\{10,3\}$, (II) $\{8,3\}$, and (III) $\{8,8\}$ lattice and varying system size $n_{l}$ at $\epsilon=0.01$.
 D.3 Quantum Monte Carlo
-Figure 17: AFM order parameter for the $\{10,3\}$, $\{8,3\}$ and $\{8,8\}$ lattice obtained by QMC calculations at $\beta t=100$ (filled symbols) and $\beta t=25$ (unfilled symbols).
-In order to supplement the picture given in Figs.
-9 and 11, we present the order parameter $m_{AFM}^{z}$ as a function of $U$ in Fig.
+Figure 18: AFM order parameter for the $\{10,3\}$, $\{8,3\}$, and $\{8,8\}$ lattice obtained by QMC calculations at $\beta t=100$ (filled symbols) and $\beta t=25$ (unfilled symbols).
+To supplement the picture given in Figs.
+10 and 12, we present the order parameter $m_{\mathrm{AFM}}^{z}$ as a function of $U$ in Fig.
 18 for two different temperatures $\beta t=25$ and $\beta t=100$.
@@
 This lattice shows a strong sensitivity to temperature already in the weak coupling regime, as can be seen in Fig.
-17(c).
+18(c).
 As discussed before, this effect is also present in the results from a mean-field analysis [Fig.
-4(a)].
+5(a)].